### v1

# On the Brun–Titchmarsh theorem

Abstract.
The classical Brun–Titchmarsh theorem gives an upper bound, which is of correct order of magnitude in the full range, for the number of primes $p\leqslant x$ satisfying $p\equiv a\,(\operatorname{mod}{q})$. We strength this inequality for $\log q/\log x$ in different ranges, improving upon previous works by Motohashi, Goldfeld, Iwaniec, Friedlander and Iwaniec, and Maynard for general or special moduli. In particular, we obtain a Burgess-like constant in the full range $q<x^{1/2-}.$ The proof is based on various estimates for character and exponential sums, by appealing to arithmetic exponent pairs and bilinear forms with algebraic trace functions from $\ell$-adic cohomology, and sums of Kloosterman sums from spectral theory of automorphic forms, as well as large value theorem for Dirichlet polynomials due to Huxley.
Key words and phrases:
Brun–Titchmarsh theorem, linear sieve, arithmetic exponent pairs, trace functions of $\ell$-adic sheaves, Kloosterman sums, character sums
2020 Mathematics Subject Classification:
11N13, 11N36, 11T23, 11L05, 11L07, 11M06, 11N75
1. Backgrounds and main results
1.1. Main results
This paper aims to understand the upper bound for the counting function
$\pi(x;q,a):=|\{p\leqslant x:p\equiv a\,(\operatorname{mod}{q})\}|,$
where $q\in\mathbf{Z}^{+}$, $(a,q)=1$ and $x\geqslant 3$. It is widely expected that
(1.1) $\pi(x;q,a)=(1+o(1))\frac{1}{\varphi(q)}\frac{x}{\log x}$
as $q,x\rightarrow+\infty$ in a very broad range, and the classical Siegel–Walfisz theorem asserts that this is valid for all $q\leqslant(\log x)^{A}$ with large but fixed $A$. No asymptotic formula is known for larger $q$ compared to $x$. Note that the Generalized Riemann Hypothesis may guarantee (1.1) for all $q\leqslant x^{1/2-\varepsilon}$.
With the aid of Brun’s sieve, Titchmarsh [Ti30] proved that
(1.2) $\max_{(a,q)=1}\pi(x;q,a)\ll\frac{x}{\varphi(q)\log{(x/q)}}$
for all $1\leqslant q<x,$ and this bears the so-called Brun–Titchmarsh theorem. In what follows, we assume $q\sim x^{\varpi}$ with $\varpi\in[0,1[$, and explore a positive constant $C(\varpi)$ as small as possible, such that
(1.3) $\max_{(a,q)=1}\pi(x;q,a)\leqslant(C(\varpi)+\varepsilon)\frac{1}{\varphi(q)}\frac{x}{\log x}$
holds for any $\varepsilon>0$ and $x>x_{0}(\varepsilon).$ Note that (1.2) guarantees the existence of $C(\varpi)$ for all $\varpi\in[0,1[$.
In this paper, we establish several choices of $C(\varpi)$ for general moduli and smooth moduli $q\sim x^{\varpi}$. We now quickly state our results.
Theorem 1.1.
Denote by $\vartheta$ the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$ Then we may take
(1.4) $C(\varpi)=\frac{16}{8-(3+2\vartheta)\varpi}$
[…]
(1.5) $\sum_{n\leqslant N}\chi(n)\ll N^{1-\frac{1}{r}}q^{\frac{r+1}{4r^{2}}+\varepsilon}$
holds for any non-trivial Dirichlet character modulo $q$ and $r\in\{1,2,3\}$. Taking $r=2$, (1.5) implies the so-called Burgess bound $L(\frac{1}{2},\chi)\ll q^{3/16+\varepsilon}$ for Dirichlet $L$-functions. The restriction $\varpi<9/20$ is quite essential in Iwaniec’s approach, and Theorem 1.1 modulo the Ramanujan–Petersson conjecture extends the validity of Iwaniec’s constant in a much wider range. Different from [Iw82], our approach relies heavily on the well-factorization of the linear sieve weight constructed by Iwaniec [Iw80] and Kloostermania developed by Deshouillers and Iwaniec [DI82]. In the prime moduli case, the above choice for $C(\varpi)$ is independent from the Ramanujan–Petersson conjecture, and this will follow from another different approach, by appealing to a recent work of Blomer, Humphries, Khan and Milinovich [BHKM20] on twisted fourth moment of Dirichlet $L$-functions to prime moduli. To echo the above convention of Burgess-like constant, we should mention that the bound in [BHKM20] implies the above Burgess bound for $L(\frac{1}{2},\chi)$ from an approach completely different from character sums; see Proposition 1.1 below.
Note that Iwaniec [Iw82] is in fact able to take $C(\varpi)=6/(3-\varpi)$ for $\varpi<9/20,$ which corresponds to $r=3$ in (1.5).
For $\varpi\geqslant 1/2$, we have the following choices for $C(\varpi)$.
Theorem 1.2.
[…]
$C(\varpi)=\begin{cases}8/(5-5\varpi),\quad&\varpi\in[1/2,~{}12/23[,\\ 32/(32-43\varpi),\quad&\varpi\in[12/23,~{}32/61[,\\ 24/(16-17\varpi),&\varpi\in[32/61,~{}8/15[,\\ 48/(40-49\varpi),&\varpi\in[8/15,~{}7/13[,\\ 16/(11-12\varpi),&\varpi\in[7/13,~{}6/11[,\\ 32/(28-35\varpi),&\varpi\in[6/11,~{}4/7[.\end{cases}$
Remark 1.
We may relax $q$ to be squarefree by using the work of Löffel [Lo16], as an extension of Lemma 4.3. It is also possible to generalize Theorem 1.2 to arbitrary moduli along with the same line.
To see the strength of Theorems 1.1 and 1.2, we would like to recall some previous admissible values of $C(\varpi)$ obtained by a series of big names.
$\varpi$ Admissible $C(\varpi)$ Authors
]0,1[]0,1[] 0 , 1 [ existence Titchmarsh [Ti30]
]0,1[]0,1[] 0 , 1 [ $2/(1-\varpi)$ van Lint & Richert [LR65]
]0,1/3[]0,1/3[] 0 , 1 / 3 [ $16/(8-3\varpi)$ Motohashi [Mo73]
$[1/3,2/5]$ $4/(2-\varpi)$ Motohashi [Mo74]
]2/5,1/$2]$ / 5 , 1 / 2 ] $2/(2-3\varpi)$ Motohashi [Mo73]
]0,24/71[]0,24/71[] 0 , 24 / 71 [ $16/(8-3\varpi)$ Goldfeld [Go75]
]0,9/20[]0,9/20[] 0 , 9 / 20 [ $16/(8-3\varpi)$ Iwaniec [Iw82, Theorem 3]
]9/20,2/3]]9/20,2/3]] 9 / 20 , 2 / 3 ] $8/(6-7\varpi)$ Iwaniec [Iw82, Theorem 6]
[6/11,1[[6/11,1[[ 6 / 11 , 1 [ $(2-(\frac{1-\varpi}{4})^{6})/(1-\varpi)$ Friedlander & Iwaniec [FI97, Theorem 1]
$[1-\delta,1[$ $(2-c_{0}(1-\varpi)^{2})/(1-\varpi)$ Bourgain & Garaev [BG14, Theorem 4]
]0,1/8[]0,1/8[] 0 , 1 / 8 [ $2$ Maynard [Ma13]
The second part of this paper is to reduce the admissible value of $C(\varpi)$ further by assuming that $q$ has very good factorizations. For a small constant $\eta>0$, we say $q$ is $\eta$-smooth if all prime factors of $q$ do not exceed $q^{\eta}.$ The assumption on the smoothness allows us to utilize the factorization of $q$ to control the resultant algebraic exponential sums very effectively. In particular, one may go far beyond the so-called Pólya–Vinogradov barrier for short exponential and character sums, as a consequence of the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs developed by Wu and Xi [WX21]. One is referred to Section 4.2 for details.
The following theorem shows that we may go further when $\varpi$ is close to $2/3$, in particular, larger than $2/3$ in a explicit way, provided that $q$ is squarefree and smooth enough.
Theorem 1.3.
Suppose $q$ is $\eta$-smooth and squarefree. Then we may take
$C(\varpi)=\frac{160}{89-91\varpi}$
[…]
$\frac{9}{51}\leqslant\varpi\leqslant\frac{9}{11}=\frac{2}{3}+\frac{5}{33},$
provided that $\eta$ is small enough.
Theorem 1.3 is in fact a special case of the following slightly general formulation on taking $(\kappa,\lambda)=(\tfrac{1}{20},\tfrac{33}{40})$.
Theorem 1.4.
[…]
$C(\varpi)=\frac{4}{(3+\kappa-\lambda)-(3+2\kappa-\lambda)\varpi}$
[…]
$\frac{1+\kappa-\lambda}{2+2\kappa-\lambda}\leqslant\varpi\leqslant\frac{1+\kappa-\lambda}{1+2\kappa-\lambda},$
provided that $\eta$ is small enough.
Theorems 1.3 and 1.4 are stronger than van Lint and Richert [LR65] and Iwaniec [Iw82] when $\varpi$ is around $2/3$ Put
$f(\kappa,\lambda)$ $=\frac{1+\kappa-\lambda}{1+2\kappa-\lambda},$
$g_{\varpi}(\kappa,\lambda)$ $=(3+\kappa-\lambda)-(3+2\kappa-\lambda)\varpi.$
To have $f(\kappa,\lambda)$ or $g_{\varpi}(\kappa,\lambda)$ as large as possible, it requires to choose $(\kappa,\lambda)$ by applying $A$- and $B$-processes to $(0,1)$ suitably; we refer to Section 4.2 for the terminology and conventions. The optimization process can be conducted following the algorithm in [GK91, Section 5.3]. In particular, for $\varpi=2/3,$ we find
$g_{\varpi}(\kappa,\lambda)$ $=1-\frac{\kappa+\lambda}{3},$
and it suffices to minimize $\kappa+\lambda.$ This task matches exactly in determining Rankin’s constant for Riemann zeta functions. According to Rankin (see also [GK91, Section 5.4]), the minimal value for $\kappa+\lambda$ should be $0.829021\cdots$ by choosing
$(\kappa,\lambda)=ABA^{3}BA^{2}BABABA^{2}BABABA^{2}BA^{2}\cdots B(0,1),$
in which case we get $g_{\varpi}(\kappa,\lambda)=0.723659\cdots$. This yields $C(2/3)=4/g_{\varpi}(\kappa,\lambda)\approx 5.2746$ in Theorem 1.4. Note that van Lint and Richert [LR65] obtained the constant $62/(1-\varpi)=$ 6 for $\varpi=2/3.$
The following theorem turns out to be better than Theorem 1.5 when $\varpi$ is a bit smaller than $1/2$.
Theorem 1.5.
Suppose $q$ is $\eta$-smooth and squarefree. Then we may take
$C(\varpi)=\begin{cases}2,\quad&\varpi\in[1/8,~{}5/12[,\\ 5/(5-6\varpi),&\varpi\in[5/12,~{}9/20[\end{cases}$
in ($,$ provided that $\eta$ is small enough.
Theorem 1.5 can be compared with the following conditional result due to Motohashi [Mo74]. Under the Linderlöf Hypothesis for Dirichlet $L$-functions, he proved that take
$C(\varpi)=\begin{cases}2,\quad&\varpi\in[0,~{}1/3],\\ 2/(2-3\varpi),&\varpi\in[1/3,~{}1/2],\end{cases}$
is admissible in (1.3).
The following two figures present our new choices of $C(\varpi)$ for $\varpi<1/2$ and $\varpi>1/2,$ respectively, and one may see when we can succeed or not.
Figure 1. $C(\varpi)$ for $\varpi<1/2$
Figure 2. $C(\varpi)$ for $\varpi>1/2$
1.2. Conditional results
As a (weaker) substitute of the Lindelöf Hypothesis, it is a reasonable approach to consider moments of Dirichlet $L$-functions. It has been known for a long time that
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll q^{\varepsilon}(1+|t|)^{O(1)}$
for any $\varepsilon>0$ and $t\in\mathbf{R}$. In addition, we may also expect, for an arbitrary coefficient $\boldsymbol{\alpha}=(\alpha_{m})$ and any $\varepsilon>0$, that
(1.6) $\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m\leqslant q^{\delta}}\alpha_{m}\chi(m)\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll(q(1+|t|))^{\varepsilon}q^{\delta}\|\boldsymbol{\alpha}\|_{\infty}$
with $\delta>0$ as large as possible. This motivates us to make the following conjecture.
Conjecture 1.1.
[…]
By extending Motohashi’s identity to link fourth moment of Dirichlet $L$-functions and cubic moment of $\mathrm{GL}_{2}$ $L$-functions, Blomer, Humphries, Khan and Milinovich [BHKM20] proved the following bound.
Proposition 1.1.
[…]
The following two theorems show that Conjecture 1.1 may guarantee sharp choices of $C(\varpi)$ for $\varpi<1/2$, by combining the possible good factorizations of $q$.
Theorem 1.6.
[…]
$C(\varpi)=\begin{cases}2,\quad&\varpi\in[1/8,~{}10/(24-5\delta)[,\\ 20/(20-(24-5\delta)\varpi),&\varpi\in[10/(24-5\delta),~{}9/20[,\end{cases}$
provided that $\eta$ is small enough.
Remark 2.
Theorem 1.5 corresponds to the degenerate case $\delta=0$ of Theorem 1.6.
Theorem 1.7.
Assume that Conjecture 1.1 holds for all $q\in\mathbf{Z}^{+}$ and some $\delta\in[0,16/45[.$ Then we may take
$C(\varpi)=\frac{4}{2-(1-\delta)\varpi}$
for $\varpi\in[9/20,1/2].$
1.3. Ingredients of proof and comparisons with previous works
Before discussing the ingredients of our proof, it would be better to give a short survey on the methods employed in previous works.
  • Both of van Lint and Richert [LR65] and Motohashi [Mo74, Mo73] utilized Selberg’s sieve, which is more convenient than Brun’s sieve at that time. Recall that one takes $C(\varpi)=2/(1-\varpi)$ in [LR65], and the constant $2$ is resulted naturally by applications of the linear sieve with its best possible main term. On the other hand, Montgomery and Vaughan [MV73] developed the large sieve, from which they deduced that $\pi(x;q,a)<2x/(\varphi(q)\log(q/x))$ for all $1\leqslant q<x$ and $(a,q)=1$. The above denominator $(1-\varpi)$ comes from the level $(x/q)^{1-\varepsilon}$ after a trivial treatment of error terms. It was Motohashi [Mo74, Mo73] who first dealt with the error term in a non-trivial manner, and he successfully employed the bilinear structure in Selberg’s sieve. He employed multiplicative characters, and a basic tool is the mean value of Dirichlet $L$-functions, which prevents him from improving van Lint and Richert [LR65] for $\varpi>1/2.$
  • Iwaniec [Iw82] developed a new linear sieve of combinatorial type with well-factorable weights. This far-reaching work was originally motivated by Chen [Ch75] and Motohashi [Mo74], and the merit is that one may optimize the balances between convolutions in the error term, and the sign changes among them can be captured more effectively. In particular, this well-factorization in the error term allows Iwaniec [Iw82] to employ (Weil’s bound for) Kloosterman sums to beat the barrier $\varpi\approx 1/2$. He was able to beat van Lint and Richert [LR65] even if $\varpi$ is close to $2/3-.$
  • For $\varpi\geqslant 2/3,$ the above approach of Iwaniec [Iw82] is no longer valid since the incomplete Kloosterman sums are extremely short, and Friedlander and Iwaniec [FI97] found another surprising approach to beat the record in [LR65]. They utilized an estimate of Karatsuba [Ka95] on the estimates for the bilinear form
    (1.7) $\sum_{m}\sum_{n}\alpha_{m}\beta_{n}\mathrm{e}\Big{(}\frac{h\overline{mn}}{q}\Big{)},$
    where $m$ and $n$, as prime variables, are restricted to very special but short segments. In particular, they succeed in improving van Lint and Richert [LR65] for all $\varpi\in[2/3,1[.$ To apply Karatsuba’s estimate for (1.7), Friedlander and Iwaniec [FI97] introduced Buchstab’s iteration so that one may create prime variables in the sifting functions. By employing ideas from additive combinatorics (sum-product phenomenon), Bourgain and Garaev [BG14] can bound (1.7) without restricting $m,n$ to primes. Therefore, Buchstab’s iteration is not necessary, and they can give a further improvement over [FI97].
  • All above works are built on sieve methods (Brun sieve, Selberg sieve, or Rosser–Iwaniec sieve). The approach of Maynard [Ma13] is purely analytic (zero density estimates, Landau–Siegel zeros, etc), and he proved that $\pi(x;q,a)<2\mathrm{Li}(x)/\varphi(q)$ for all $c<q\leqslant x^{1/8},$ where $\mathrm{Li}(x)=\int_{2}^{x}\mathrm{d}t/\log t$ and $c$ is a computable constant. Moreover, he is also able to produce lower bounds for $\pi(x;q,a)$ with effectively computable constants.
Our work improves the above results in several aspects for different values of $\varpi.$ In principle, we need to bound the difference of convolutions
$\mathop{\sum_{m\sim M}\sum_{n\sim N}\sum_{l\sim L}}_{mnl\equiv a\,(\operatorname{mod}{q})}\alpha_{m}\beta_{n}-\frac{1}{\varphi(q)}\mathop{\sum_{m\sim M}\sum_{n\sim N}\sum_{l\sim L}}_{(mnl,q)=1}\alpha_{m}\beta_{n}$
with $MN$ as large as possible, where $MNL=x.$ The coefficient $-β\boldsymbol{\alpha}*\bbeta$ comes from the linear sieve of Iwaniec [Iw82], and each of $-β\boldsymbol{\alpha},\bbeta$ is bounded by divisor functions. There are at least two approaches to detect the congruence condition $mnl\equiv a\,(\operatorname{mod}{q}),$ via multiplicative (Dirichlet) characters or additive characters.
Multiplicative Approach: After a direct application of orthogonality of multiplicative characters, the problem might reduce to proving
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m\sim M}\sum_{n\sim N}\sum_{l\sim L}\alpha_{m}\beta_{n}\chi(mnl)\Big{|}\ll x(\log x)^{-2024}.$
One may easily see this approach does not work unless $\varpi<1/2,$ by expecting that there exist (at most) squareroot cancellations on the left hand side. After suitable applications of Hölder’s inequality, one may arrive at bounding the twisted moment
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{r\sim R}c_{r}\chi(r)\Big{|}^{2}\Big{|}\sum_{l\sim L}\chi(l)\Big{|}^{2k}.$
The underlying idea is to prove squareroot cancellations among the averages over $r$ and $l$. This expectation cannot be realized when $R$ or $k$ is suitably large, hence it becomes necessary to control large values of Dirichlet polynomials.
In his original proof, Iwaniec [Iw82] applied fourth moment of short character sums or Dirichlet $L$-functions (i.e., $k=2$ and $R=1$). As an approximation to the sixth moment, we may also allow, in the case $k=2$, the length $R$ to be suitably large so that a kind of Linderlöf Hypothesis on average can be achieved. This will appeal to the recent work of Blomer, Humphries, Khan and Milinovich [BHKM20], in which $q$ is restricted to primes and $R$ can be as large as $q^{1/4}$.
Note that $(c_{r})$ has a convolution structure, coming from Iwaniec’s new construction of the Rosser–Iwaniec sieve. We utilize the special structure of $(c_{r})$ as efficiently as possible, and for $k=2$ we may give a another kind of Linderlöf Hypothesis on average for $R\approx q^{(5-2\vartheta)/8}$. Together with large sieves (Lemma 3.1) and large value estimate for Dirichlet polynomials (Lemma 3.2), this can also produce a level of distribution in linear sieves with the same strength as the above approach. The advantage is that $q$ can be an arbitrary large integer, but we are restricted by the possibility of exceptional spectrum of the non-Eucliean Laplacian. Modulo the Ramanujan–Petersson conjecture, we reach exactly the same level.
Additive Approach: A direct application of Poisson summation will reduce the problem to proving
$\frac{1}{H}\sum_{h\leqslant H}\mathop{\sum_{m\sim M}\sum_{n\sim N}}_{(mn,q)=1}\alpha_{m}\beta_{n}\mathrm{e}\Big{(}\frac{h\overline{mn}}{q}\Big{)}\ll MN(\log x)^{-2024},$
where $H\approx qMN/x.$ It is natural to apply Cauchy–Schwarz to smooth out one variable, say $m,$ and we then arrive at the incomplete Kloosterman sum
$\sum_{\begin{subarray}{c}m\sim M\\ (m,q)=1\end{subarray}}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
The Pólya–Vinogradov method transforms the incomplete sum to a complete one, and then Weil’s bound for individual Kloosterman sums applies. This was exactly done by Iwaniec [Iw82]. Our approach differs from Iwaniec’s in two directions:
(1) If $q$ is squarefree and has only very small prime factors, the factorization of $q$ can be controlled very well, so that the work on arithmetic exponent pairs, developed by Wu and Xi [WX21], can be applied. In such situation, one can save much more than before given the incomplete Kloosterman sum of the same length, and one can also estimate the incomplete sum non-trivially even though $M$ is smaller than $q^{1/2}$. In particular, this approach works very well when $\varpi$ is around $2/3$. The details can be found in Section 6.5.
(2) If $q$ has no good factorizations, say $q$ is a large prime, we also employ the Pólya–Vinogradov method, but do not go with Weil’s bound directly. Alternatively, we take the advantages of averages over $h_{1},h_{2},n_{1},n_{2}$, as well as the dual sum resulted from the Pólya–Vinogradov method. By grouping $h_{1},h_{2},n_{1},n_{2}$ in certain ways, we may create some bilinear and trilinear forms with Kloosterman sums. This produces possibilities to capture sign changes of Kloosterman sums. For the bilinear form, we need to control a sum of product of many Kloosterman sums, which will appeal to [FGKM14] or [Xi17]; while for the trilinear form, we reduce the problem to bounding the bilinear form
$\sum_{m}\sum_{n}K(mn),$
where $K:\mathbf{Z}/q\mathbf{Z}\rightarrow\mathbf{C}$ is given by the Fourier transform of a product of two Kloosterman sums. Hence the work of Fouvry, Kowalski and Michel [FKM14] on smooth bilinear forms with general trace functions will apply, which is rooted in $\ell$-adic cohomology and spectral theory of Eisenstein series. The details can be found in the first several parts of Section 6, as well as Section 4.1.
1.4. Connections with Landau–Siegel zeros
It should be mentioned that the reduction of $C(\varpi)$ is of great importance for the location of Landau–Siegel zeros. Motohashi [Mo79] showed that if for $x>q^{c},$
$\pi(x;q,a)<(2-\delta)\frac{1}{\varphi(q)}\frac{x}{\log(x/q)}$
holds with some constant $\delta>0$, the real zero $\beta$ (if exists) of $L(s,\chi)$ then satisfies
$\beta<1-\frac{c^{\prime}\eta}{\log(q+3)},$
implying there are no Landau–Siegel zeros. The phenomenon that such improvements over the Brun-Titchmarsh theorem eliminate Landau–Siegel zeros seems first observed by Klimov [Kl61, Remark 1]; see also [Si83] and [Gr22] for related results.
On the other hand, it is believed that $\pi(x,a;q)\leqslant 2\mathrm{Li}(x)/\varphi(q)$ as long as $q$ is at most a small power of $x$; see [Mo12] for the details and history. Recently, Maynard [Ma13] demonstrated this expectation is valid for all $q^{8}\leqslant x.$ His method involves a log-free zero density estimate for Dirichlet $L$-functions and careful analysis of Landau–Siegel zeros. Up to a very small constant $\varepsilon$, we have shown in Theorem 1.5 that $\pi(x,a;q)$ is at most twice as much as the conjectural bound, provided that $q<x^{5/12},$ $q$ is squarefree and all prime factors are sufficiently small.
Notation and convention
We write $\mathrm{e}(t)=\mathrm{e}^{2\pi\mathrm{i}t}$. We use $\varepsilon$ to denote a small positive number, which might be different at each occurrence. The notation $n\sim N$ means $N<n\leqslant 2N.$ We write $A\leqslant B^{-}$ if $A\leqslant B^{1-\delta}$ holds for some constant $\delta>0.$
Acknowledgements
This work is supported in part by NSFC (No. 12025106).
2. Linear sieves
We introduce some conventions and fundamental results on the linear Rosser–Iwaniec sieve. Let $\mathcal{A}=(a_{n})$ be a finite sequence of integers and $\mathcal{P}$ a set of prime numbers. Define the sifting function
$S(\mathcal{A},\mathcal{P},z)=\sum_{(n,P(z))=1}a_{n}\quad\text{with}\quad P(z)=\prod_{p<z,\,p\in\mathcal{P}}p.$
For squarefree $d$ with all its prime factors belonging to $\mathcal{P}$, we consider the subsequence $\mathcal{A}_{d}=(a_{n})_{n\equiv 0\!\,(\operatorname{mod}{d})}$ and the congruence sum
$A_{d}=\sum_{n\equiv 0\!\,(\operatorname{mod}{d})}a_{n}.$
In applications to sieve methods, it is expected that $\mathcal{A}$ equidistributes in the special arithmetic progression $n\equiv 0\,(\operatorname{mod}{d})$: There are an appropriate approximation $\mathfrak{M}$ to $A_{1}$ and a multiplicative function $g$ supported on squarefree numbers with all its prime factors belonging to $\mathcal{P}$ verifying
$0<g(p)<1\quad(p\in\mathcal{P})$
such that
(a) the remainder
$r(\mathcal{A},d)=A_{d}-g(d)\mathfrak{M}$
is small on average over $d\mid P(z)$;
(b) there exists a constant $L>1$ such that
$\frac{V(z_{1})}{V(z_{2})}\leqslant\frac{\log z_{2}}{\log z_{1}}\bigg{(}1+\frac{L}{\log z_{1}}\bigg{)}\quad\text{with}\quad V(z)=\prod_{p<z,\,p\in\mathcal{P}}(1-g(p))$
for $2\leqslant z_{1}<z_{2}$.
Let $F$ and $f$ be the continuous solutions to the system
(2.1) $\begin{cases}sF(s)=2\mathrm{e}^{\gamma}&\text{for $\,0<s\leqslant 2$},\\ sf(s)=0&\text{for $\,0<s\leqslant 2$},\\ (sF(s))^{\prime}=f(s-1)&\text{for $\,s>2$},\\ (sf(s))^{\prime}=F(s-1)&\text{for $\,s>2$},\end{cases}$
where $\gamma$ is the Euler constant.
For $k\geqslant 1$, denote by $\tau_{k}(n)$ the number of ways of expressing $n$ as the product of $k$ positive integers. An arithmetic function $\lambda(d)$ is of level $D$ and order $k$, if
$\lambda(d)=0\quad(d>D)\qquad\text{and}\qquad|\lambda(d)|\leqslant\tau_{k}(d)\quad(d\leqslant D).$
Let $r\geqslant 2$ be a positive integer. We say that $\lambda$ is well-factorable of degree $r$, if for every decomposition $D=D_{1}D_{2}\cdots D_{r}$ with $D_{1},D_{2},\dots,D_{r}\geqslant 1,$ there exist $r$ arithmetic functions $\lambda_{1},\lambda_{2},\dots,\lambda_{r}$ such that
$\lambda=\lambda_{1}*\lambda_{2}*\cdots*\lambda_{r}$
with each $\lambda_{j}$ of level $D_{j}$ and order $k$.
We now state the following fundamental result of Iwaniec [Iw80].
Lemma 2.1.
Let $r\geqslant 2$ and $0<\varepsilon<\tfrac{1}{8}$. Under the above hypothesis, we have
$S(\mathcal{A},\mathcal{P},z)\leqslant\mathfrak{M}V(z)\bigg{\{}F\bigg{(}\frac{\log D}{\log z}\bigg{)}+E\bigg{\}}+\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}(d)r(\mathcal{A},d)$
for all $z\geqslant 2$, where $F$ is given by the system (2.1), $T$ depends only on $\varepsilon$, $\lambda_{t}(d)$ is well-factorable of level $D,$ degree $r$ and order $1$, and $E\ll\varepsilon+\varepsilon^{-8}\mathrm{e}^{L}(\log D)^{-1/3}.$
As one can show, $F(s)$ decays rapidly to $1$ as $s\rightarrow+\infty.$ In applications, we would like to explore the admissible $D$ as large as possible such that the total remainder
$\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}(d)r_{d}(\mathcal{A})$
is still under control. In principle, a larger $D$ would produce a better upper bound for $S(\mathcal{A},\mathcal{P},z)$ upon a suitable choice for $z.$ The merit of Lemma 2.1 lies in the well-factorization of $\lambda_{t}$, so that one can estimate the resultant bilinear forms more effectively to produce a better admissible $D$ (compared with earlier sieve methods before Iwaniec). This bilinear structure is very crucial in many existing applications of Iwaniec’s linear sieve, and in fact only bilinear structure is employed in such applications. As one may see from Lemma 2.1, we can take $r$ to be a large integer, so that some multilinear structure appears, and a natural question is whether we can utilize this new structure effectively in practice. As an instance of this feature, we need to express $\lambda_{t}$, in the proof of Theorem 1.1 for prime moduli, as a convolution of four arithmetic functions.
3. Dirichlet polynomials and character sums
The following lemma follows from the orthogonality of Dirichlet characters; see [Mo71, Theorem 6.2].
Lemma 3.1.
For an arbitrary coefficient $\boldsymbol{\alpha}=(\alpha_{n})_{n\sim N},$ we have
$\sum_{\chi\,(\operatorname{mod}{q})}\Big{|}\sum_{n\sim N}\alpha_{n}\chi(n)\Big{|}^{2}\ll(N+q)\|\boldsymbol{\alpha}\|^{2}.$
We also need Huxley’s large value theorem for Dirichlet polynomials; see [Hu74] for the original proof, as well as Jutila [Ju77] for another elegant proof.
Lemma 3.2.
For an arbitrary coefficient $\boldsymbol{\alpha}=(\alpha_{n})_{n\sim N}$ and $V>0,$ we have
#$\#\Big{\{}\chi\,(\operatorname{mod}{q}):\Big{|}\sum_{n\sim N}\alpha_{n}\chi(n)\Big{|}>V\Big{\}}\ll\|\boldsymbol{\alpha}\|^{2}NV^{-2}+\|\boldsymbol{\alpha}\|^{6}q^{1+\varepsilon}NV^{-6}.$
In the proof of Theorem 1.1, we may reduce the problem to bounding the second moment of Dirichlet $L$-functions weighted by a Dirichlet polynomial:
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m}c_{m}\chi(m)\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{2}.$
Alternatively, we may also consider the second moment of incomplete character sums weighted by a Dirichlet polynomial:
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m}c_{m}\chi(m)\Big{|}^{2}\Big{|}\sum_{n\leqslant N}\chi(n)n^{it}\Big{|}^{2}.$
We hope to prove the Lindelöf Hypothesis on average with the support of $c_{m}$ as large as possible. In what follows, we turn to a slightly general situation.
Let $F\in\mathcal{C}_{0}^{\infty}(\mathbf{R}^{2})$ be a Schwartz function with support in $]N,2N]\times]N,2N]$, satisfying
$\frac{\partial^{\nu_{1}+\nu_{2}}}{\partial x_{1}^{\nu_{1}}\partial x_{2}^{\nu_{2}}}F(x_{1},x_{2})$ $\ll\eta^{\nu_{1}+\nu_{2}}N^{-\nu_{1}-\nu_{2}}$
[…]
We have the following asymptotic formula.
Theorem 3.1.
[…]
(3.2) $\mathcal{R}$ $=\frac{\varphi(q)}{q^{2}}\Big{(}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M\\ (m_{1}m_{2},q)=1\end{subarray}}\alpha_{m_{1}}\beta_{m_{2}}\Big{)}\Big{(}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z\Big{)}+\mathcal{E},$
where
$\mathcal{E}\ll\eta^{\frac{9}{2}}M^{\frac{3}{2}}Nq^{\frac{\vartheta-1}{2}}(1+Mq^{-\frac{1}{2}})^{\frac{3}{2}}(MNq)^{\varepsilon}+q^{-1}(MN)^{2}(\log MNq)^{-A}+\eta(MN)^{1+\varepsilon}$
for any $\varepsilon,A>0,$ with an implied constant depending only on $(\varepsilon,A)$. Here $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
The control on the error term in Theorem 3.1 depends heavily on well-factorizations of $-β\boldsymbol{\alpha},\bbeta.$ The proof will be given in Section 7, and we appeal to Kloostermania in the following version.
Lemma 3.3.
Let $C,M,N\geqslant 1,$ $a,r,s\in\mathbf{Z}^{+}$ pairwise coprime, and $g(c,m,n)$ a smooth function with compact support in $[C,2C]\times[M,2M]\times[N,2N]$ such that
$\frac{\partial^{\nu_{1}+\nu_{2}+\nu_{3}}}{\partial c^{\nu_{1}}\partial m^{\nu_{2}}\partial n^{\nu_{3}}}g(c,m,n)\ll c^{-\nu_{1}}m^{-\nu_{2}}n^{-\nu_{3}}$
for all $\boldsymbol{\nu}=(\nu_{1},\nu_{2},\nu_{3})\in\mathbf{N}^{3},$ the implied constant in $\ll$ depending at most on $\boldsymbol{\nu}.$
For an arbitrary complex coefficient $\boldsymbol{\beta}=(\beta_{n}),$ we have
$\mathop{\sum_{c}\sum_{m}\sum_{n}}_{(sc,rm)=1}\beta_{n}g(c,m,n)\mathrm{e}\Big{(}\frac{\pm an\overline{rm}}{sc}\Big{)}$ $\ll(arsCMN)^{\varepsilon}\|\boldsymbol{\beta}\|\cdot\Delta,$
where
$\Delta$ $=a^{\vartheta}(rsCM)^{\frac{1}{2}}\frac{(1+1/X)^{\vartheta}}{1+X^{2}}\Big{(}1+X+\frac{C}{rM}\Big{)}^{\frac{1}{2}}\Big{(}1+X+\frac{N}{rs}\Big{)}^{\frac{1}{2}}+s^{-1}MN^{\frac{1}{2}}$
with $X=aN/(srCM).$ Here $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
Lemma 3.3 is essentially due to Deshouillers and Iwaniec [DI82], and the above version is directly taken from Matomäki [Ma09, Lemma 9].
Theorem 3.1 allows us to derive the following version of Lindelöf Hypothesis on average.
Lemma 3.4.
Let $q\geqslant 3$ and $g$ a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=(\lambda_{n})_{n\leqslant N}$ be a well-factorable function of level $N.$ Then we have
$\frac{1}{\varphi(q)}$ $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n}\lambda_{n}\chi(n)\Big{|}^{2}\Big{|}\sum_{l\in\mathbf{Z}}g\Big{(}\frac{l}{L}\Big{)}\chi(l)\Big{|}^{2}$
$\ll\eta^{\frac{9}{2}}N^{\frac{3}{2}}Lq^{\frac{\vartheta-1}{2}}(1+Nq^{-\frac{1}{2}})^{\frac{3}{2}}q^{\varepsilon}+\Big{(}\eta^{\frac{9}{2}}+\frac{NL}{q}\Big{)}NL(\log q)^{-A}$
for any $\varepsilon,A>0,$ where the implied constant depends only on $(\varepsilon,A).$
The following lemma is an unconditional case of Conjecture 1.1, provided that the Dirichlet polynomial is not too long.
Lemma 3.5.
Let $q$ be a large prime and $N\leqslant q^{\frac{1}{4}}$. Let $\boldsymbol{\alpha}=(\alpha_{n})_{n\leqslant N}$ be an arbitrary coefficient supported on squarefree numbers. Then we have
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n\leqslant N}\alpha_{n}\chi(n)\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll(1+|t|)^{12}Nq^{\varepsilon}\|\boldsymbol{\alpha}\|_{\infty}$
for any $\varepsilon>0$ and $t\in\mathbf{R},$ where the implied constant depends only on $\varepsilon.$
As stated in Proposition 1.1, Lemma 3.4 is proved by Blomer, Humphries, Khan and Milinovich [BHKM20] for $t=0$, and their argument can be easily extended to all $t\in\mathbf{R}$. It is very useful to relax $q$ to an arbitrary modulus, which in turn can guarantee Theorem 1.1 with $\vartheta=0$ for all $q\geqslant 3$.
4. General trace functions and arithmetic exponent pairs
Parts of this paper will employ $\ell$-adic cohomology from algebraic geometry to study the necessary exponential and character sums. Many of them would be formulated in the language of trace functions which will be recalled below.
4.1. Trace functions over finite fields
For an odd prime $p$, we consider the middle-extension sheaves on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$. Recall that a sheaf $\mathcal{F}$ on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$ is said to be middle-extension, if $\mathcal{F}\simeq j_{*}(j^{*}\mathcal{F})$ holds for each $j:U\rightarrow\mathbf{A}^{1}$ as the open immersion of a non-empty open set on which $\mathcal{F}$ is lisse. Fix an isomorphism $\iota:\overline{\mathbf{Q}}_{\ell}\rightarrow\mathbf{C}$ with another prime $\ell\neq p$. Following Katz [Ka88, Section 7.3.7], we define trace functions as follows.
Definition 4.1 (Trace functions).
Let $\mathcal{F}$ be an $\ell$-adic middle-extension sheaf pure of weight zero, which is lisse on an open set $U$. The trace function associated to $\mathcal{F}$ is defined by
$K:x\in\mathbf{F}_{p}\mapsto\iota(\mathrm{tr}(\mathrm{Frob}_{x}\mid V_{\mathcal{F}})),$
where $\mathrm{Frob}_{x}$ denotes the geometric Frobenius at $x\in\mathbf{F}_{p},$ and $V_{\mathcal{F}}$ is a finite dimensional $\overline{\mathbf{Q}}_{\ell}$-vector space, which is corresponding to a continuous finite-dimensional Galois representation and unramified at every closed point $x$ of $U.$
Motivated by many applications in analytic number theory, the conductor of a (middle-extension) sheaf was introduced by Fouvry, Kowalski and Michel [FKM14] to measure the geometric complexity of the trace function. For an $\ell$-adic middle-extension sheaf $\mathcal{F}$ on $\mathbf{P}^{1}_{\mathbf{F}_{p}}$ of rank $\mathrm{rank}(\mathcal{F})$, the conductor of $\mathcal{F}$ is defined to be
$\mathfrak{c}(\mathcal{F})=\mathrm{rank}(\mathcal{F})+\sum_{x\in S(\mathcal{F})}(1+\mathrm{Swan}_{x}(\mathcal{F})),$
where $S(\mathcal{F})\subset\mathbf{P}^{1}(\overline{\mathbf{F}}_{p})$ denotes the ((((finite)))) set of singularities of $\mathcal{F}$, and $\mathrm{Swan}_{x}(\mathcal{F})$ $(\geqslant 0)$ denotes the Swan conductor of $\mathcal{F}$ at $x$.
We have a lot of practical examples of trace functions arising in modern analytic number theory. Among many interesting examples, we would like to mention additive characters, multiplicative characters, and Kloosterman sums.
  • Let $f\in\mathbf{F}_{p}(X)$ be a rational function, and $\psi$ a primitive additive character on $\mathbf{F}_{p}$, then $\psi(f(x))$ is a trace function of an $\ell$-adic middle-extension sheaf, which is taken to be zero when meeting a pole of $f$ at $x$. More precisely, one can show that there exists an $\ell$-adic middle-extension sheaf modulo $p$, denoted by $\mathcal{L}_{\psi(f)},$ such that $x\mapsto\psi(f(x))$ is the trace function of $\mathcal{L}_{\psi(f)}.$ The conductor can be bounded in terms of the degree of $f$, independent of $p$.
  • Let $f\in\mathbf{F}_{p}(X)$ be a rational function, and $\chi$ a multiplicative character of order $d>1$. If $f$ has no pole or zero of order divisible by $d$, then one can show that there exists an $\ell$-adic middle-extension sheaf, denoted by $\mathcal{L}_{\chi(f)}$, such that $x\mapsto\chi(f(x))$ is the trace function of $\mathcal{L}_{\chi(f)}$. The conductor can also be bounded in terms of the degree of $f$, independent of $p$.
  • Another example is the following normalized Kloosterman sum defined by
    $\mathrm{Kl}(\cdot,p):x\in\mathbf{F}_{p}\mapsto\frac{1}{\sqrt{p}}\sum_{y\in\mathbf{F}_{p}^{\times}}\mathrm{e}\Big{(}\frac{y+x/y}{p}\Big{)}.$
    According to Deligne, there exists an $\ell$-adic middle-extension sheaf $\mathcal{K}\ell$, called a Kloosterman sheaf, such that $\mathrm{Kl}(\cdot,p)$ gives its trace function over $\mathbf{F}_{p}^{\times}.$ The sheaf $\mathcal{K}\ell$ is geometrically irreducible and is of rank $2$, the conductor of which is bounded by $5$.
An middle-extension sheaf on $\mathbf{P}_{\mathbf{F}_{p}}^{1}$, which is pointwise pure of weight 0 is said to be a Fourier sheaf if no geometrically irreducible component is geometrically isomorphic to an Artin–Schreier sheaf $\mathcal{L}_{\psi}$ attached to some additive character $\psi$ of $\mathbf{F}_{p}.$
The following lemma collects some properties of Fourier transforms of Fourier sheaves due to Deligne [De80], Laumon [La87], Brylinski [Br86], Katz [Ka88, Ka90] and Fouvry, Kowalski and Michel [FKM15].
Lemma 4.1.
Let $\psi$ be a non-trivial additive character of $\mathbf{F}_{p}$ and $\mathcal{F}$ a Fourier sheaf on $\mathbf{P}_{\mathbf{F}_{p}}^{1}$ with the trace function $K_{\mathcal{F}}.$ Then there exists an $\ell$-adic sheaf $\mathcal{G}=\mathrm{FT}_{\psi}(\mathcal{F})$ called the Fourier transform of $\mathcal{F}$, which is also an $\ell$-adic Fourier sheaf with the trace function
$K_{\mathcal{G}}(y)=\mathrm{FT}_{\psi}(K_{\mathcal{F}})(y):=-\frac{1}{\sqrt{p}}\sum_{x\in\mathbf{F}_{p}}K_{\mathcal{F}}(x)\psi(yx).$
Furthermore, we have
  • The sheaf $\mathcal{G}$ is geometrically irreducible, or geometrically isotypic, if and only if $\mathcal{F}$ is;
  • The Fourier transform is involutive, in the sense that we have a canonical arithmetic isomorphism $\mathrm{FT}_{\psi}(\mathcal{G})\simeq[\times(-1)]^{*}\mathcal{F},$ where [×$[\times(-1)]^{*}$ denotes the pull-back by the map $x\mapsto-x;$
  • We have $\mathfrak{c}(\mathrm{FT}_{\psi}(\mathcal{F}))\leqslant 10\mathfrak{c}(\mathcal{F})^{2}.$
For $a,b,y\in\mathbf{F}_{p}$, put
$V_{p}(y;a,b)=\frac{1}{\sqrt{p}}\sum_{x\in\mathbf{F}_{p}}\mathrm{Kl}(ax,p)\mathrm{Kl}(bx,p)\mathrm{e}\Big{(}\frac{-yx}{p}\Big{)}.$
In the proof of Theorem 1.2, we need to show that $y\mapsto V_{p}(y;a,b)$ is a good trace function in some sense, provided that $a\neq b$.
Lemma 4.2.
Let $\mathcal{K}\ell$ be the Kloosterman sheaf of rank $2$ as above. For all $a,b\in\mathbf{F}_{p}^{\times}$ with $a\neq b,$ the tensor product :=([×$\mathcal{F}:=([\times a]^{*}\mathcal{K}\ell)\otimes([\times b]^{*}\mathcal{K}\ell)$ is a Fourier sheaf, which is geometrically irreducible, pure of weight zero, of rank $4,$ and has the geometric monodromy group $\mathrm{SL}_{2}\times\mathrm{SL}_{2}.$ The conductor of $\mathcal{F}$ is at most $25$
Moreover, for each non-trivial additive character $\psi$ of $\mathbf{F}_{p},$ the Fourier transform $\mathrm{FT}_{\psi}(\mathcal{F})$ is also Fourier sheaf, which is geometrically irreducible and of conductor at most $6250$
Proof.
The first part has been proven by Fouvry and Michel [FM07, Section 7]. The second part follows from Lemma 4.1 immediately. ∎
Before going into more details on trace functions over finite fields, we would like mention another definition of Kloosterman sums defined over $\mathbf{Z}/q\mathbf{Z}:$ for $q\geqslant 1$, put
$S(m,n;q)=\sideset{}{{}^{*}}{\sum}_{a\,(\operatorname{mod}{q})}\mathrm{e}\Big{(}\frac{ma+n\overline{a}}{q}\Big{)},$
where $m,n\in\mathbf{Z}$ and $a\overline{a}\equiv 1\,(\operatorname{mod}{q}).$ It is clear that $\mathrm{Kl}(x,p)\sqrt{p}=S(x,1;p)=S(1,x;p)$ for each prime $p$, if we identify $\mathbf{F}_{p}$ by $\{0,1,\cdots,p-1\}.$ We say $S(*,0;q)$ a Ramanujan sum, which satisfies $|S(m,0;q)|\leqslant(m,q);$ in general, we have the Weil bound
$|S(m,n;q)|\leqslant q^{\frac{1}{2}}(m,n,q)^{\frac{1}{2}}\tau(q).$
We now recall a very strong estimate of Fouvry, Kowalski and Michel [FKM14] on smooth bilinear forms with general trace functions.
Lemma 4.3.
Let $K$ be a trace function of an $\ell$-adic middle-extension sheaf $\mathcal{F}$ on $\mathbf{A}_{\mathbf{F}_{p}}^{1},$ which is geometrically isotypic, pointwise pure of weight $0,$ and no geometrically irreducible component of which is geometrically isomorphic to an Artin–Schreier sheaf $\mathcal{L}_{\psi}$ attached to some additive character $\psi$ of $\mathbf{F}_{p}.$
Let $M,N\geqslant 1$ and $V_{1},V_{2}$ two fixed smooth functions with compact supports in $[1,2].$ We then have
$\mathop{\sum\sum}_{m,n\in\mathbf{Z}}K(mn)V_{1}\Big{(}\frac{m}{M}\Big{)}V_{2}\Big{(}\frac{n}{N}\Big{)}\ll MN\Big{(}1+\frac{p}{MN}\Big{)}^{\frac{1}{2}}p^{-\frac{1}{8}+\varepsilon}$
for any $\varepsilon>0,$ where the implicit constant depends only on $\varepsilon$ and the conductor of $\mathcal{F}.$
Lemma 4.4.
[…]
$\sum_{m\,(\operatorname{mod}{p})}\Big{|}\sum_{n\in\mathcal{N}}\beta_{n}\mathrm{Kl}(mn,p)\Big{|}^{2\nu}\ll|\mathcal{N}|^{\nu}p+|\mathcal{N}|^{2\nu}p^{\frac{1}{2}},$
where the implied constant depends only on $\nu.$
Proof.
Denote by $S$ the quantity in question. We may write
$|S|\leqslant\sum_{\mathbf{n}\in\mathcal{N}^{2\nu}}|S(\mathbf{n})|,$
where for $\mathbf{n}=(n_{1},n_{2},\ldots,n_{2\nu})\in\mathcal{N}^{2\nu}$,
$S(\mathbf{n})=\sum_{x\,(\operatorname{mod}{p})}\prod_{1\leqslant j\leqslant 2\nu}\mathrm{Kl}(n_{j}x).$
There is a trivial bound that
$S(\mathbf{n})\ll p.$
We also expect there exist squareroot cancellations if the coordinates of $\mathbf{n}$ appear in suitable configurations. More precisely, according to [FGKM14] or [Xi17], we have
$S(\mathbf{n})\ll\sqrt{p}$
as long as the multiplicity
$\mu_{n}(\mathbf{n}):=|\{1\leqslant j\leqslant 2\nu:n_{j}=n\}|$
is odd for some $n\in\mathcal{N}$. In this way, we employ the above trivial bound if $\mu_{n}(\mathbf{n})$ is even for all $n\in\mathcal{N}$. This proves the lemma readily. ∎
4.2. Composite trace functions and arithmetic exponent pairs
By the Chinese remainder theorem, a trace function can be also defined for squarefree moduli, which allows Wu and Xi [WX21] to develop the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs.
In terms of trace functions over finite fields, we may construct composite trace functions following the manners in [Po14] and [WX21]. Let $q\geqslant 1$ be a squarefree number and $K_{q}$ a composite trace function (formally) defined by the product
(4.1) $K_{q}(n)=\prod_{p\mid q}K_{p}(n),$
where $K_{p}$ is the trace function of certain $\ell$-adic middle-extension sheaf $\mathcal{F}_{p}$ on $\mathbf{A}^{1}_{\mathbf{F}_{p}}$ for each $p\mid q$. We adopt the convention that $K_{q}(n)=1$ for all $n$ if $q=1.$ In practice, the value of $K_{p}(n)$ may depend on the complementary divisor $q/p.$ Originally, $K_{p}$ is a function living in $\mathbf{Z}/p\mathbf{Z}$; one can extend its domain to $\mathbf{Z}$ by periodicity, and one can do the same thing to $K_{q}$.
For a given interval $I$, we consider the following average
$\sum_{n\in I}K_{q}(n).$
Roughly speaking, the Pólya–Vinogradov method can be used to bound $\mathfrak{S}(K,W)$ non-trivially for $|I|>q^{\frac{1}{2}+\varepsilon}$, and it was observed by Heath-Brown [HB78] that one may go further if $q$ has food factorizations, and this led him to prove a Weyl-type subconvexity bound for $L(\tfrac{1}{2},\chi)$ to smooth moduli. Inspired by Heath-Brown, Wu and Xi [WX21] developed the method of arithmetic exponent pairs for averages of composite trace functions that may go far beyond the Pólya–Vinogradov bound, as long as the factorization of $q$ is good enough. We now summarize results of the arithmetic exponent pairs developed in [WX21]. For more backgrounds and applications of the $q$-analogue of van der Corput method and arithmetic exponent pairs, the readers may refer to [WX21] directly and references therein.
Following Sawin [WX21], we call a middle-extension sheaf $\mathcal{F}_{p}$ on $\mathbf{A}_{\mathbf{F}_{p}}^{1}$ universally amiable if
  1. (i)
    it is pointwise pure of weight 00;
  2. (ii)
    no geometrically irreducible component of $\mathcal{F}_{p}$ is geometrically isomorphic to an Artin–Schreier sheaf $\mathcal{L}_{\psi}$ attached to some additive characer $\psi$ of $\mathbf{F}_{p}$;
  3. (iii)
    its local monodromy at $\infty$ has all slopes $\leqslant 1$.
In such case, we also say the associated trace function $K_{p}$ is universally-amiable. A composite trace function $K_{q}$ is said to be compositely universally amiable if for each $p\mid q,$ $K_{p}$ can be decomposed into a sum of universally amiable trace functions, in which case we also say the corresponding sheaf $\mathcal{F}:=(\mathcal{F}_{p})_{p\mid q}$ is compositely universally amiable.
To mention the main results in [WX21], we assume that
  • $\eta>0$ is a sufficiently small number;
  • $q$ is a squarefree number with no prime factors exceeding $q^{\eta}$;
  • $\delta$ is positive integer such that $(\delta,q)=1$;
  • $K_{q}$ is a compositely universally amiable trace function $\,(\operatorname{mod}{q})$;
  • $I$ is an interval with $|I|<q\delta;$
  • there exists some uniform constant $\mathfrak{c}>0,$ such that $\mathfrak{c}(\mathcal{F}_{p})\leqslant\mathfrak{c}$ for each $p\mid q$.
We further assume that $\delta$ is also squarefree and no prime factors exceed $\delta^{\eta}$, as will be satisfied in our applications in this paper. We are interested in the upper bound of the average
$\mathfrak{S}(K,W)=\sum_{n\in I}K_{q}(n)W_{\delta}(n),$
where $W_{\delta}:\mathbf{Z}/\delta\mathbf{Z}\rightarrow\mathbf{C}$ is an arbitrary function, which we call deformation factor later. Let $(\kappa,\lambda,\nu)$ be a tuple such that
(4.2) $\mathfrak{S}(K,W)\ll_{\eta,\varepsilon,\mathfrak{c}}N^{\varepsilon+O(\eta)}\|W_{\delta}\|_{\infty}(q/N)^{\kappa}N^{\lambda}\delta^{\nu}$
for any $\varepsilon>0,$ where the implied constant is allowed to depend on $\eta,\varepsilon$ and $\mathfrak{c}.$ We define two maps
(4.3) $A\cdot(\kappa,\lambda,\nu)=\Big{(}\frac{\kappa}{2(\kappa+1)},~{}\frac{\kappa+\lambda+1}{2(\kappa+1)},~{}\frac{\kappa}{2(\kappa+1)}\Big{)}$
and
(4.4) $B\cdot(\kappa,\lambda,\nu)=\Big{(}\lambda-\frac{1}{2},~{}\kappa+\frac{1}{2},~{}\lambda+\nu-\kappa\Big{)},$
resulted from $A$- and $B$-processes in the $q$-analogue of van der Corput method; see [WX21] for details. Note that (4.2) holds trivially for $(\kappa,\lambda,\nu)=(0,1,0).$ Under different combinations of $A$ and $B$, we may produce a series of tuples starting from $(\kappa,\lambda,\nu)=(0,1,0).$ Namely, we may introduce
(4.5) $\mathfrak{P}=\{A^{m_{1}}B^{n_{1}}A^{m_{2}}B^{n_{2}}\cdots A^{m_{k}}B^{n_{k}}\cdot(0,1,0):k\geqslant 1,m_{j}\geqslant 0,n_{j}\in\{0,1\}\}.$
By induction, one may show that
[…]
for each $(\kappa,\lambda,\nu)\in\mathfrak{P}.$ In practice, the exponent $\nu$ is not quite crucial (since it is at most $1$) and we formally call the elements in $\mathfrak{P}$ as arithmetic exponent pairs. Note that this convention is different from that in [WX21].
We are now ready to formulate the work in [WX21] on arithmetic exponent pairs.
Lemma 4.5.
[…]
$(\kappa,\lambda,\nu)\in\mathfrak{P}.$
Amongst many choices, one may take $(\kappa,\lambda,\nu)$ to be
$B\cdot(0,1,0)$ $=\Big{(}\frac{1}{2},\frac{1}{2},1\Big{)},$
$AB\cdot(0,1,0)$ $=\Big{(}\frac{1}{6},\frac{2}{3},\frac{1}{6}\Big{)},$
$BA^{2}B\cdot(0,1,0)$ $=\Big{(}\frac{2}{7},\frac{4}{7},\frac{11}{14}\Big{)},$
$A^{2}BA^{2}B\cdot(0,1,0)$ $=\Big{(}\frac{1}{20},\frac{33}{40},\frac{1}{20}\Big{)},$
or
$A^{k}B\cdot(0,1,0)$ $=\Big{(}\frac{1}{2^{k+1}-2},1-\frac{k}{2^{k+1}-2},\frac{1}{2^{k+1}-2}\Big{)}$
for each positive integer $k\geqslant 2$.
As a direct consequence of Lemma 4.5, we have the following estimate for incomplete character sums to smooth moduli.
Lemma 4.6.
Let $\eta>0$ be a sufficiently small number. Let $q$ be an $\eta$-smooth squarefree number and $\chi$ a non-trivial Dirichlet character modulo $q.$
For each positive integer $k$ and an interval $I,$ we have
$\sum_{n\in I}\chi(n)\ll q^{\kappa}|I|^{\lambda-\kappa}q^{O(\eta)}$
with
$(\kappa,\lambda)=\Big{(}\frac{1}{2^{k+1}-2},1-\frac{k}{2^{k+1}-2}\Big{)},$
where the implied constant depends only on $(\eta,k).$ Consequently, for each interval $I$ with $|I|\sim q^{\theta}$ and $\theta\in]0,1],$ there exists some constant $c=c(\theta)>0$ such that
$\sum_{n\in I}\chi(n)\ll|I|^{1-c+O(\eta)},$
where the implied constant depends only on $(\eta,\theta).$
Proof.
The first part follows directly from Lemma 4.5. We then have
$\sum_{n\in I}\chi(n)\ll|I|^{1-c+O(\eta)}$
with
$c=\frac{1}{2^{k+1}-2}\Big{(}k-1-\frac{1}{\theta}\Big{)}.$
Taking $k=[3/\theta]$ finishes the proof. ∎
The following lemma gives an upper bound for incomplete Kloosterman sums to smooth moduli.
Lemma 4.7.
Let $\eta>0$ be a sufficiently small number. Let $q$ be an $\eta$-smooth squarefree number. For each integer $h$ and an interval $I,$ we have
$\sum_{\begin{subarray}{c}n\in I\\ (n,q)=1\end{subarray}}\mathrm{e}\Big{(}\frac{h\overline{n}}{q}\Big{)}\ll\frac{|I|}{q}(h,q)+q^{\kappa}|I|^{\lambda-\kappa+{O(\eta)}}(h,q)^{\nu},$
[…]
Proof.
We first assume that $|I|<q$. Let $d=(h,q),q_{1}=q/d$ and $h_{1}=h/d$. Hence
$S:=\sum_{\begin{subarray}{c}n\in I\\ (n,q)=1\end{subarray}}\mathrm{e}\Big{(}\frac{h\overline{n}}{q}\Big{)}$ $=\sum_{\begin{subarray}{c}n\in I\\ (n,q_{1})=1\end{subarray}}\mathrm{e}\Big{(}\frac{h_{1}\overline{n}}{q_{1}}\Big{)}\mathbf{1}_{(n,d)=1}.$
We are now in a good position to apply Lemma 4.5 with
$(q,\delta)\leftarrow(q_{1},d),\ \ K_{q_{1}}:x\mapsto\mathrm{e}\Big{(}\frac{h_{1}\overline{x}}{q_{1}}\Big{)},\ \ W_{d}:x\mapsto\mathbf{1}_{(x,d)=1},$
getting
$S\ll q^{\kappa}|I|^{\lambda-\kappa+{O(\eta)}}(h,q)^{\nu}.$
For $|I|\geqslant q,$ the average $S$ would be split into $O(|I|/q)$ copies of Ramanujan sum $S(h,0,q)$ and one incomplete sum of length at most $q$. This incomplete sum can be bounded as above, and the lemma is proven by noting that $|S(h,0,q)|\leqslant(h,q)$. ∎
5. Proofs of Theorems 1.1, 1.5
5.1. Initial treatment by linear sieves
Let $y=x(\log x)^{-3}$ and introduce a smooth function $\varPhi$ which is supported on $[y,x+y]$, satisfying
$\begin{cases}\varPhi(t)=1\ \ &\text{for }t\in[2y,x],\\ \varPhi(t)\geqslant 0\ \ &\text{for }t\in\mathbf{R},\\ \varPhi^{(j)}(t)\ll_{j}y^{-j}&\text{for all }j\geqslant 0,\end{cases}$
and $\widehat{\varPhi}(0)=\int_{\mathbf{R}}\varPhi(t)\mathrm{d}t=x.$ Therefore,
$\pi(x;q,a)=\pi^{*}(x;q,a)+O(q^{-1}x(\log x)^{-2}),$
where
$\pi^{*}(x;q,a)=\sum_{p\equiv a\,(\operatorname{mod}{q})}\varPhi(p).$
In view of the definition of $C(\varpi)$ in (1.3), the above error $O(q^{-1}x(\log x)^{-2})$ is acceptable, and it suffices to derive upper bounds for the smooth sum $\pi^{*}(x;q,a).$
For $(d,q)=1$ we consider the congruence sum
(5.1) $A(d):=\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ d\mid n\end{subarray}}\varPhi(n)=\sum_{n\equiv a\overline{d}\,(\operatorname{mod}{q})}\varPhi(dn).$
Define
(5.2) $r(d)$ $:=A(d)-\frac{x}{dq}.$
Let $(\lambda_{d})$ be a linear upper bound sieve of level $D$, so that $1*\mu\leqslant 1*\lambda$. Then
$\pi^{*}(x;q,a)\leqslant\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{d}A(d)\leqslant\frac{x}{q}V(z)F\Big{(}\frac{\log D}{\log z}\Big{)}+\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{d}r(d),$
where $F$ is defined by (2.1) and
$V(z):=\prod_{p<z,\,p\nmid q}\Big{(}1-\frac{1}{p}\Big{)}=\frac{q}{\varphi(q)}\cdot\frac{\mathrm{e}^{-\gamma}}{\log z}\Big{\{}1+O\Big{(}\frac{1}{\log x}\Big{)}\Big{\}}.$
Thanks to Lemma 2.1, we may choose well-factorable remainder terms in the above application of linear Rosser–Iwaniec sieve. We thus conclude the following result.
Proposition 5.1.
Let $\varepsilon>0$ and $D<x^{1-\varepsilon}.$ We have
(5.3) $\pi(x;q,a)\leqslant\{2+O(\varepsilon)\}\frac{x}{\varphi(q)\log D}+\sum_{t\leqslant T(\varepsilon)}\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{t}(d)r(d),$
[…]
(5.4) $R(D):=\sum_{d\leqslant D}\lambda(d)r(d)\ll\frac{x}{q(\log x)^{2024}}$
with $D$ as large as possible.
5.2. Approaching by multiplicative characters
From orthogonality of Dirichlet characters, it follows that
$r(d)$ $=\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\sum_{n\in\mathbf{Z}}\varPhi(dn)\chi(dn)+O(1).$
By Mellin inversion, we may write
$r(d)$ $=\frac{1}{\varphi(q)}\frac{1}{2\pi i}\int_{(2)}\widetilde{\varPhi}(s)d^{-s}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\chi(d)L(s,\chi)\mathrm{d}s+O(1),$
where $L(s,\chi)$ is the Dirichlet $L$-function. In view of the decay
$\widetilde{\varPhi}(s)$ $\ll x^{\Re(s)}(1+x^{-1}y|s|)^{-A}$
for any $A\geqslant 0$, we shift the integral to $\Re(s)=1/2$ and sum over $d\leqslant D,$ getting
$R(D)$ $=\frac{1}{\varphi(q)}\frac{1}{2\pi}\int_{\mathbf{R}}\widetilde{\varPhi}(\tfrac{1}{2}+it)\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\mathcal{D}(\tfrac{1}{2}+it,\chi)L(\tfrac{1}{2}+it,\chi)\mathrm{d}t+O(1),$
where $\mathcal{D}(s,\chi)$ is a Dirichlet polynomial defined by
$\mathcal{D}(s,\chi)=\sum_{d\leqslant D}\lambda(d)\chi(d)d^{-s}.$
Note that
$\int_{\mathbf{R}}|\widetilde{\varPhi}(\tfrac{1}{2}+it)|(1+|t|)^{3}\mathrm{d}t$ $\ll x^{\frac{1}{2}}(\log x)^{12},$
it suffices to prove that for $\Re(s)=\frac{1}{2},$
(5.5) $\Big{|}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\mathcal{D}(s,\chi)L(s,\chi)\Big{|}\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
Due to the well-factorization of $\lambda$, we may write
[…]
where $\boldsymbol{\alpha}$, $-β\bbeta$ and $\boldsymbol{\gamma}$ are of level $M$, $N$ and $W$ respectively, and
[…]
$\mathcal{M}(s,\chi)=\sum_{m\sim M}\alpha_{m}\chi(m)m^{-s},\ \ \mathcal{N}(s,\chi)=\sum_{n\sim N}\beta_{n}\chi(n)n^{-s},\ \ \mathcal{W}(s,\chi)=\sum_{w\sim W}\gamma_{w}\chi(w)w^{-s}.$
Hence the proof (5.4) can be reduced to proving that
(5.7) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)L(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}$
for $\Re(s)=\frac{1}{2}.$ Recall that $\varPhi$ is supported in $[y,x+y]$. For $m\sim M,n\sim N,w\sim W,$ it is clear that $\varPhi(mnwl)$ vanishes unless $l\asymp x/(MNW)$. As argued above, an alternative approach to (5.4) is proving
(5.8) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036},$
where $\Re(s)=\frac{1}{2},$ and
$\mathcal{L}(s,\chi)=\sum_{l\in\mathbf{Z}}\phi\Big{(}\frac{l}{L}\Big{)}\chi(l)l^{-s},\ \ \ L=\frac{x}{MNW}.$
Here $\phi$ is a fixed smooth function which dominates the indicator function of $[-1,1].$
5.3. Concluding Theorem 1.1
We assume that $W=1$ in (5.6), so that (5.8) reduces to
(5.9) $\Xi:=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
By Cauchy–Schwarz inequality, we have
$\Xi^{2}\ll(M+q)\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|^{2}.$
From Lemma 3.4 it follows that
$\Xi^{2}\ll|s|^{\frac{9}{2}}(M+q)N^{\frac{1}{2}}q^{\frac{\vartheta+1}{2}+\varepsilon}(1+Nq^{-\frac{1}{2}})^{\frac{3}{2}}+(M+q)(|s|^{\frac{9}{2}}q+xM^{-1}(\log q)^{-A})$
for any $A>0$, provided $NL>q^{1+\varepsilon}.$ It is easy to check that (5.9) is valid as long as
$M=x/q^{1+\varepsilon},\ \ N=q^{\frac{5-2\vartheta}{8}-\varepsilon}$
if $q<x^{\frac{1}{2}-}$, in which case we may take $D=x/q^{\frac{3+2\vartheta}{8}+2\varepsilon}$ in Proposition 5.1. This proves the first part of Theorem 1.1 readily.
We now assume that $q$ is a large prime. As argued above, (5.7) reduces to
(5.10) $\Theta:=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)L(s,\chi)|\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
By Hölder’s inequality, we have
(5.11) $\Theta^{4}\ll\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)|^{\frac{4}{3}}\Big{)}^{3}\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)L(s,\chi)|^{4}\Big{)}.$
If $N^{2}\leqslant q^{\frac{1}{4}}$, we infer from Lemma 3.5 that
(5.12) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{N}(s,\chi)L(s,\chi)|^{4}\ll|s|^{12}q^{1+\varepsilon}.$
To bound the $\frac{4}{3}$-th moment of $\mathcal{M}(s,\chi),$ we assume that $\boldsymbol{\alpha}=\boldsymbol{\alpha}_{1}*\boldsymbol{\alpha}_{2}*\boldsymbol{\alpha}_{3}$ is a triple convolution of three coefficients $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}$ and $\boldsymbol{\alpha}_{3},$ each of which is of level $M^{\frac{1}{3}}$. This is reasonable since $\boldsymbol{\alpha}$ comes from the well-factorable weight constructed by Iwaniec [Iw82]; see Lemma 2.1. Note that $\mathcal{M}_{j}(s,\chi)$ denotes the Dirichlet polynomial generated by $\boldsymbol{\alpha}_{j}$. We now derive from Lemma 3.1 that
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}(s,\chi)|^{\frac{4}{3}}$ $=\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}_{1}(s,\chi)\mathcal{M}_{2}(s,\chi)\mathcal{M}_{3}(s,\chi)|^{\frac{4}{3}}$
$\ll\prod_{1\leqslant j\leqslant 3}\Big{(}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{M}_{j}(s,\chi)|^{4}\Big{)}^{\frac{1}{3}}$
$\ll(M^{\frac{2}{3}}M^{\varepsilon}+q).$
from which and (5.12), (5.11), we find (5.10) is valid as long as
$M=x/q^{\frac{1}{2}+2\varepsilon},\ \ N=q^{\frac{1}{8}},$
if $q<x^{\frac{1}{2}}(\log x)^{-2036}$, in which case we may take $D=x/q^{\frac{3}{8}+2\varepsilon}$ in Proposition 5.1. The second part of Theorem 1.1 is established.
By the way, if assuming Conjecture 1.1, we may take $N=q^{\delta/2}$, so that the choice $D=x/q^{\frac{1-\delta}{2}+2\varepsilon}$ is admissible in Proposition 5.1. This proves Theorem 1.7.
5.4. Concluding Theorem 1.5
In this subsection, we assume that $q$ is squarefree and $\eta$-smooth, as required in Theorem 1.5. The application of Lemma 3.2 will depend on the sizes of corresponding Dirichlet polynomials, we divide the non-trivial characters $\,(\operatorname{mod}{q})$ into eight sets. Precisely, we define
$\Gamma_{1}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{1}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{2}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{2}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{3}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{3}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{4}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{4}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\}.$
For $1\leqslant k\leqslant 4$, and $s=\frac{1}{2}+it$ with $t\in\mathbf{R},$ we define
$S_{k}^{\pm}$ $=\sum_{\begin{subarray}{c}\chi\in\Gamma_{k}^{\pm}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)L(s,\chi)|,$
$T_{k}^{\pm}$ $=\sum_{\begin{subarray}{c}\chi\in\Gamma_{k}^{\pm}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|.$
[…]
(5.13) $\min\{S_{k}^{*},~{}T_{k}^{*}\}\ll|s|^{3}x^{\frac{1}{2}}(\log x)^{-2036}.$
To estimate these sums, we first develop some lemmas on mean values of $\mathcal{M}(s,\chi),$ $\mathcal{N}(s,\chi),$ $\mathcal{L}(s,\chi)$ and $L(s,\chi)$ for $s=\frac{1}{2}+it,$ which would be denoted by $\mathcal{M},\mathcal{N},\mathcal{L}$ and $L$ for convenience.
The following lemma is a direct consequence of Lemma 3.1.
Lemma 5.1.
For $1\leqslant k\leqslant 4$, we have
$\sum_{\chi\in\Gamma_{k}^{\pm}}|\mathcal{M}|^{2}\ll M+q$
and
$\sum_{\chi\in\Gamma_{k}^{\pm}}|\mathcal{N}|^{2}\ll N+q.$
Lemma 5.2.
For $\Gamma\in\{\Gamma_{1}^{+},\Gamma_{1}^{-},\Gamma_{2}^{+},\Gamma_{2}^{-}\},$ we have
(5.14) $\sum_{\chi\in\Gamma}|\mathcal{M}|^{2}\ll Mq^{\varepsilon}.$
For $\Gamma\in\{\Gamma_{1}^{+},\Gamma_{1}^{-},\Gamma_{3}^{+},\Gamma_{3}^{-}\},$ we have
(5.15) $\sum_{\chi\in\Gamma}|\mathcal{N}|^{2}\ll Nq^{\varepsilon}.$
For $\Gamma\in\{\Gamma_{3}^{+},\Gamma_{3}^{-},\Gamma_{4}^{+},\Gamma_{4}^{-}\},$ we have
(5.16) $\sum_{\chi\in\Gamma}|\mathcal{M}|^{6}\ll Mq^{1+\varepsilon}.$
For $\Gamma\in\{\Gamma_{2}^{+},\Gamma_{4}^{-},\Gamma_{4}^{+},\Gamma_{4}^{-}\},$ we have
(5.17) $\sum_{\chi\in\Gamma}|\mathcal{N}|^{6}\ll Nq^{1+\varepsilon}.$
Proof.
For $\Gamma\in\{\Gamma_{1}^{+},\Gamma_{1}^{-},\Gamma_{2}^{+},\Gamma_{2}^{-}\}$, Lemma 3.2 yields
$\sum_{\chi\in\Gamma}|\mathcal{M}|^{2}$ $\ll\log x\mathop{\sup}\limits_{q^{\frac{1}{4}}\leqslant A\ll M^{\frac{1}{2}}}A^{2}(MA^{-2}+q^{1+\varepsilon}MA^{-6})\ll Mq^{\varepsilon}.$
The proofs of other inequalities are similar. ∎
Following similar arguments, we also have following estimates.
Lemma 5.3.
For $1\leqslant k\leqslant 4,$ we have
(5.18) $\sum_{\chi\in\Gamma_{k}^{+}}|\mathcal{W}\mathcal{L}|^{4}\ll(WL)^{2}q^{\varepsilon}$
(5.19) $\sum_{\chi\in\Gamma_{k}^{-}}|\mathcal{W}\mathcal{L}|^{12}\ll(WL)^{2}q^{1+\varepsilon}.$
We are now ready to prove (5.13). Without loss of generality, we assume $q>x^{0.001}.$ Fix $\varepsilon>0$ to be a very small constant in terms of $\eta$. We also assume $L\geqslant q^{\varepsilon}$, so that
(5.20) $MNW\leqslant x^{1-\varepsilon}.$
From Lemma 4.6, it follows that
(5.21) $\mathcal{L}(s,\chi)\ll|s|L^{\frac{1}{2}-}.$
We also assume that
(5.22) $W\leqslant q^{\frac{\delta}{2}},$
so that Conjecture 1.1 yields
(5.23) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{W}(s,\chi)L(s,\chi)|^{4}\ll q^{1+\varepsilon}.$
(I) Treatment to $T_{1}^{\pm}$: From (5.21) and the trivial bound $\mathcal{W}\ll W^{\frac{1}{2}}$, we obtain
$T_{1}^{\pm}\ll|s|(WL)^{\frac{1}{2}-}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{M}||\mathcal{N}|.$
By Cauchy–Schwarz inequality and Lemma 5.1, we have
$T_{1}^{\pm}\ll|s|(WL)^{\frac{1}{2}-}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{2}}\ll|s|x^{\frac{1}{2}-}$
as desired.
(II) Treatments to $S_{2}^{\pm}$ and $T_{2}^{\pm}$: As argued above, we have
$T_{2}^{\pm}\ll|s|x^{\frac{1}{2}-}$
for $N\geqslant q$. We henceforth assume that $N<q.$ Using Hölder’s inequality,
$S_{2}^{\pm}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\chi\in\Gamma_{2}^{\pm}}|\mathcal{W}L|^{4}\Big{)}^{\frac{1}{4}}.$
Hence Lemma 5.1, (5.17) and (5.23) yield
$S_{2}^{\pm}\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{1}{8}}(Nq)^{\frac{1}{8}}q^{\frac{1}{4}+\varepsilon}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}.$
Using (5.18) or (5.19) in place of (5.23), we obtain
$T_{2}^{+}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\chi\in\Gamma_{2}^{+}}|\mathcal{W}\mathcal{L}|^{4}\Big{)}^{\frac{1}{4}}$
$\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{1}{8}}(Nq)^{\frac{1}{8}}(W^{2}L^{2})^{\frac{1}{4}}$
$\ll q^{\frac{1}{4}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}(WL)^{\frac{1}{2}}$
and
$T_{2}^{-}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{3}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{24}}\Big{(}\sum_{\chi\in\Gamma_{2}^{-}}|\mathcal{W}\mathcal{L}|^{12}\Big{)}^{\frac{1}{12}}$
$\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{3}{8}}(Nq)^{\frac{1}{24}}(qW^{2}L^{2})^{\frac{1}{12}}$
$\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{24}}(WL)^{\frac{1}{6}}.$
Collecting the above bounds, we have
$\min\{S_{2}^{+},~{}T_{2}^{+}\}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}\min\{1,q^{-\frac{1}{4}}(WL)^{\frac{1}{2}}\}\ll x^{\frac{1}{8}}q^{\frac{7}{16}}M^{\frac{3}{8}}q^{\varepsilon}$
and
$\min\{S_{2}^{-},~{}T_{2}^{-}\}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}\min\{1,N^{-\frac{1}{12}}(WL)^{\frac{1}{6}}\}\ll x^{\frac{1}{12}}q^{\frac{1}{2}}M^{\frac{5}{12}}q^{\varepsilon}.$
[…]
(5.24) $M\leqslant x^{1-\varepsilon}q^{-\frac{6}{5}},\ \ q\leqslant x^{\frac{5}{11}-\varepsilon}.$
[…]
(5.25) $N\leqslant x^{1-\varepsilon}q^{-\frac{6}{5}},\ \ q\leqslant x^{\frac{5}{11}-\varepsilon}.$
(IV) Treatments to $S_{4}^{\pm}$ and $T_{4}^{\pm}$: Firstly, we may follow the above arguments to derive that
$T_{4}^{\pm}\ll|s|x^{\frac{1}{2}-}$
for $M,N\geqslant q$. The case $N<q<M$ can be treated along with the arguments in (II), and by interchanging $M$ and $N$ one can also deal with the case $M<q<N.$ Hence we also need the restrictions (5.24) and (5.25).
We henceforth assume that $M,N<q$, in which case Hölder’s inequality gives
$S_{4}^{\pm}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\chi\in\Gamma_{4}^{+}}|\mathcal{W}L|^{4}\Big{)}^{\frac{1}{4}}.$
From Lemma 5.1, (5.16), (5.17) and (5.23), it follows that
$S_{4}^{\pm}\ll(\log x)^{2}q^{\frac{5}{16}}q^{\frac{5}{16}}(Mq)^{\frac{1}{16}}(Nq)^{\frac{1}{16}}q^{\frac{1}{4}+\varepsilon}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}.$
Using (5.18) or (5.19) in place of (5.23), we obtain
$T_{4}^{+}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{W}\mathcal{L}|^{4}\Big{)}^{\frac{1}{4}}$
$\ll(\log x)^{2}q^{\frac{5}{16}}q^{\frac{5}{16}}(Mq)^{\frac{1}{6}}(Nq)^{\frac{1}{6}}(W^{2}L^{2})^{\frac{1}{4}}$
$\ll q^{\frac{3}{4}+\varepsilon}(MN)^{\frac{1}{16}}(WL)^{\frac{1}{2}}$
and
$T_{4}^{-}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{7}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{7}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{48}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{48}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{W}\mathcal{L}|^{12}\Big{)}^{\frac{1}{12}}$
$\ll(\log x)^{2}q^{\frac{7}{16}}q^{\frac{7}{16}}(Mq)^{\frac{1}{48}}(Nq)^{\frac{1}{48}}(qW^{2}L^{2})^{\frac{1}{12}}$
$\ll q^{1+\varepsilon}(MN)^{\frac{1}{48}}(WL)^{\frac{1}{6}}.$
Collecting the above bounds, we have
$\min\{S_{4}^{+},~{}T_{4}^{+}\}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}\min\{1,q^{-\frac{1}{4}}(WL)^{\frac{1}{2}}\}\ll x^{\frac{1}{16}}q^{\frac{31}{32}+\varepsilon}$
and
$\min\{S_{4}^{-},~{}T_{4}^{-}\}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}\min\{1,(WL)^{\frac{1}{6}}(MN)^{-\frac{1}{24}}\}\ll x^{\frac{1}{20}}q^{1+\varepsilon}.$
[…]
Combining the above treatments in four parts, as well as the restrictions (5.20), (5.24), (5.25) and (5.22), we choose
$M=N=x^{1-\varepsilon}q^{-\frac{6}{5}},\ \ W=\min\{x^{1-\varepsilon}(MN)^{-1},~{}q^{\frac{\delta}{2}}\}$
if $q\leqslant x^{\frac{9}{20}-\varepsilon}$, in which case we may take
$D=\min\{x^{1-\varepsilon},~{}x^{2-\varepsilon}q^{\frac{\delta}{2}-\frac{12}{5}}\}$
in Proposition 5.1. This proves Theorem 1.5 readily.
6. Proof of Theorems 1.2 and 1.4
Keep in mind that $q\geqslant x^{\frac{1}{2}}$ throughout this section.
6.1. Approaching by additive characters
Recall that
$A(d)=\sum_{n\equiv a\overline{d}\,(\operatorname{mod}{q})}\varPhi(dn)$
and $\widehat{\varPhi}(0)=x.$ It follows from Poisson summation that
$A(d)=\frac{1}{dq}\sum_{h\in\mathbf{Z}}\widehat{\varPhi}\Big{(}\frac{h}{dq}\Big{)}\mathrm{e}\Big{(}\frac{ah\overline{d}}{q}\Big{)}.$
From the rapid decay of $\widehat{\varPhi}$, we may write
$r(d)$ $=A(d)-\frac{x}{dq}=\frac{1}{dq}\sum_{0<|h|\leqslant H}\widehat{\varPhi}\Big{(}\frac{h}{dq}\Big{)}\mathrm{e}\Big{(}\frac{ah\overline{d}}{q}\Big{)}+O(x^{-1})$
for $H:=MNqX^{-1+\varepsilon}$.
Take $\lambda$ to be well-factorable of degree $2$; that is,
$-β\displaystyle\lambda=\boldsymbol{\alpha}*\bbeta,$
where $\boldsymbol{\alpha}$ and $-β\bbeta$ are of level $M$ and $N$, respectively, and
[…]
It suffices to show that
$\sum_{m\leqslant M}\sum_{n\leqslant N}\alpha_{m}\beta_{n}r(mn)\ll\frac{x}{q(\log x)^{2024}}.$
By dyadic device, it suffices to prove that
(6.1) $\mathcal{R}(M,N):=\sum_{m\sim M}\sum_{n\sim N}\frac{\alpha_{m}\beta_{n}}{mn}\sum_{0<|h|\leqslant H}\widehat{\varPhi}\Big{(}\frac{h}{mnq}\Big{)}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{q}\Big{)}\ll\frac{x}{(\log x)^{2026}},$
where $MN=D$ and $H=x^{-1}MNq^{1+\varepsilon}$.
6.2. Transforming exponential sums
We transform the above multiple exponential sum in $R(M,N)$ into averages of complete Kloosterman sums. To do so, we first write
$\mathcal{R}(M,N)$ $=\sum_{m\sim M}\sum_{n\sim N}\frac{\alpha_{m}\beta_{n}}{mn}\sum_{0<|h|\leqslant H}\int_{\mathbf{R}}\varPhi(t)\mathrm{e}\Big{(}\frac{-ht}{mnq}\Big{)}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{q}\Big{)}\mathrm{d}t$
$=\sum_{m\sim M}\sum_{n\sim N}\frac{\alpha_{m}\beta_{n}}{n}\sum_{0<|h|\leqslant H}\int_{\mathbf{R}}\varPhi(mt)\mathrm{e}\Big{(}\frac{-ht}{nq}\Big{)}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{q}\Big{)}\mathrm{d}t$
$\ll\frac{1}{N}\int_{\mathbf{R}}\sum_{m\sim M}\varPhi(mt)\Big{|}\sum_{0<|h|\leqslant H}\sum_{n\sim N}\beta_{h,n,t}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{q}\Big{)}\Big{|}\mathrm{d}t,$
where $|\beta_{h,n,t}|\leqslant 1.$ Therefore,
$\mathcal{R}(M,N)$ $\ll\frac{x}{MN}\sum_{m\in\mathbf{Z}}W\Big{(}\frac{m}{M}\Big{)}\Big{|}\sum_{0<|h|\leqslant H}\sum_{n\sim N}\beta_{h,n,t}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{q}\Big{)}\Big{|}$
for some $t\in\mathbf{R}$ and a smooth function $W$ which dominates the indicator function of $[1,2].$
By Cauchy–Schwarz inequality, we obtain
$\mathcal{R}(M,N)^{2}$ $\ll\frac{x^{2}}{MN^{2}}\sum_{m\in\mathbf{Z}}W\Big{(}\frac{m}{M}\Big{)}\Big{|}\sum_{0<|h|\leqslant H}\sum_{n\sim N}\beta_{h,n,t}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{q}\Big{)}\Big{|}^{2}.$
Squaring out and switching summations, it follows that
$\mathcal{R}(M,N)^{2}$ $\ll\frac{x^{2}}{MN^{2}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{\begin{subarray}{c}m\in\mathbf{Z}\\ (m,q)=1\end{subarray}}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
The diagonal terms with $h_{1}n_{2}=h_{2}n_{1}$ contribute at most
$\ll\frac{x^{2}}{N^{2}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}=h_{2}n_{1}\end{subarray}}|\beta_{h_{1},n_{1},t}\beta_{h_{2},n_{2},t}|\ll x^{2}HN^{-1}q^{\varepsilon}.$
To ensure the desired bound (6.1), we require that
(6.2) $M\leqslant q^{-1}x^{1-\varepsilon}.$
We now have
(6.3) $\mathcal{R}(M,N)^{2}$ $\ll x^{2}HN^{-1}q^{\varepsilon}+\frac{x^{2}}{MN^{2}}\varSigma$
with
(6.4) $\varSigma$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{\begin{subarray}{c}m\in\mathbf{Z}\\ (m,q)=1\end{subarray}}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
From Poisson summation, it follows that
$\varSigma$ $=\frac{M}{q}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{k\in\mathbf{Z}}\widehat{W}\Big{(}\frac{kM}{q}\Big{)}S(k,a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}};q).$
Thanks to the rapid decay of $\widehat{W},$ we may truncate the $k$-sum at $K:=M^{-1}q^{1+\varepsilon}$ with a negligible error term. Note that the zero-th frequency contributes to $\varSigma$ at most
$\ll\frac{M}{q}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\leqslant N\\ h_{1}n_{2}\neq h_{2}n_{1}\end{subarray}}(h_{1}n_{2}-h_{2}n_{1},q)\ll q^{-1+\varepsilon}M(HN)^{2}.$
Hence
(6.5) $\varSigma$ $=\varSigma^{*}+O(q^{-1+\varepsilon}M(HN)^{2})$
with
$\varSigma^{*}$ $=\frac{M}{q}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{1\leqslant|k|\leqslant K}\widehat{W}\Big{(}\frac{kM}{q}\Big{)}S(k,a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}};q).$
By employing Weil’s bound for individual Kloosterman sums, one may produce an upper bound for $\varSigma^{*}$, and thus for $\mathcal{R}(M,N).$ This was exactly done by Iwaniec [Iw82]. We would like to estimate $\varSigma^{*}$ by controlling the sign changes of Kloosterman sums effectively. More precisely, we will group certain variables and create bilinear or trilinear forms to capture cancellations among Kloosterman sums.
There exist at least two means to group variables:
  • For $u\,(\operatorname{mod}{q}),$ put
    $\rho(u)$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1},~{}(n_{1}n_{2},q)=1\\ (h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}}\equiv h\,(\operatorname{mod}{q})\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}.$
  • For $b\leqslant B$ and $c\leqslant C$ with $B=4HN$ and $C=4N^{2},$ put
    $\rho(b,c)$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1},~{}(n_{1}n_{2},q)=1\\ h_{1}n_{2}-h_{2}n_{1}=b,~{}n_{1}n_{2}=c\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}.$
Hence
$\varSigma^{*}$ $=\frac{M}{q}\sum_{u\,(\operatorname{mod}{q})}\rho(u)\sum_{1\leqslant|k|\leqslant K}\delta_{k}S(k,au;q)$
$=\frac{M}{q}\sum_{b\leqslant B}\sum_{c\leqslant C}\rho(b,c)\sum_{1\leqslant|k|\leqslant K}\delta_{k}S(k,ab\overline{c};q),$
where $\delta_{k}=\widehat{W}(kM/q).$ In the remaining of this section, we specialize $q$ to be a large prime with $M>2q^{\varepsilon}$, so that $K<q/2$ and
(6.6) $\varSigma^{*}$ $=\frac{M}{\sqrt{q}}\sum_{u\,(\operatorname{mod}{q})}\rho(u)\sum_{1\leqslant|k|\leqslant K}\delta_{k}\mathrm{Kl}(aku,q)$
(6.7) $=\frac{M}{\sqrt{q}}\sum_{b\leqslant B}\sum_{c\leqslant C}\rho(b,c)\sum_{1\leqslant|k|\leqslant K}\delta_{k}\mathrm{Kl}(akb\overline{c},q),$
where $\mathrm{Kl}(\cdot,q)=S(\cdot,1;q)/\sqrt{q}.$
6.3. Bounding $\varSigma^{*}$ as bilinear forms
For each $\nu\in\mathbf{Z}^{+},$ we infer from (6.6) and Hölder’s inequality that
$\varSigma^{*}$ $\ll\frac{M}{\sqrt{q}}\mathcal{B}_{1}^{1-\frac{1}{\nu}}(\mathcal{B}_{2}\mathcal{B}_{3})^{\frac{1}{2\nu}},$
where
$\mathcal{B}_{j}=\sum_{u\,(\operatorname{mod}{q})}\rho(u)^{j},\ \ j=1,2,$
and
$\mathcal{B}_{3}$ $=\sum_{u\,(\operatorname{mod}{q})}\Big{|}\sum_{1\leqslant|k|\leqslant K}\delta_{k}\mathrm{Kl}(ku,q)\Big{|}^{2\nu}.$
Trivially, we have
(6.8) $\mathcal{B}_{1}$ $\ll(HN)^{2}.$
From Lemma 4.4, we also find
(6.9) $\mathcal{B}_{3}$ $\ll K^{\nu}q+K^{2\nu}q^{\frac{1}{2}}.$
To bound $\mathcal{B}_{2}$, we need the following lemma.
Lemma 6.1.
With the above notation, we have
$\mathcal{B}_{2}\ll(1+H/N)(1+HN^{3}/q)(HN)^{2}q^{\varepsilon}$
for any $\varepsilon>0.$
Proof.
We first write
$\mathcal{B}_{2}$ $\ll\mathop{\sum\sum\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}\left(h_{1}n_{2}-h_{2}n_{1}\right)n_{1}^{\prime}n_{2}^{\prime}\equiv(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}\,(\operatorname{mod}{q})\\ n_{1},n_{1}^{\prime},n_{2},n_{2}^{\prime}\sim N,~{}1\leqslant|h_{1}|,|h_{1}^{\prime}|,|h_{2}|,|h_{2}^{\prime}|\leqslant H\\ (n_{1}n_{1}^{\prime}n_{2}n_{2}^{\prime},q)=1\\ (h_{1}n_{2}-h_{2}n_{1})(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})\neq 0\end{subarray}}1.$
Note that the above congruence restriction $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}\equiv(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}\,(\operatorname{mod}{q})$ is equivalent to
(6.10) $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime})n_{1}n_{2}+qt$
for some $t\in[-T,T]$ and $T:=1+4HN^{3}/q$.
Fix $n_{1}^{\prime},n_{2}^{\prime},h_{1},h_{2},t$ firstly, and choose a pair $(h_{1}^{*},h_{2}^{*})$ such that
(6.11) $(h_{1}n_{2}-h_{2}n_{1})n_{1}^{\prime}n_{2}^{\prime}=(h_{1}^{*}n_{2}^{\prime}-h_{2}^{*}n_{1}^{\prime})n_{1}n_{2}+qt,\ \ h_{1}^{*}n_{2}^{\prime}-h_{2}^{*}n_{1}^{\prime}\neq 0$
holds for some $n_{1},n_{2}.$ This may create $h_{1}^{\prime}=h_{1}^{*}+sn_{1}^{\prime}/(n_{1}^{\prime},n_{2}^{\prime})$ and $h_{2}^{\prime}=h_{2}^{*}+sn_{2}^{\prime}/(n_{1}^{\prime},n_{2}^{\prime})$ in (6.10) for some $w$ with $|s|\leqslant 2(n_{1}^{\prime},n_{2}^{\prime})H/N.$ Note that the number of such $s$ is at most $O(1+(n_{1}^{\prime},n_{2}^{\prime})H/N).$ Hence the problem reduces to counting solutions to the equation (6.11), which in fact implies
$(Cn_{1}-B)(Cn_{2}-A)=AB-Cqt,$
where
$A=-h_{1}n_{1}^{\prime}n_{2}^{\prime},\ \ B=h_{2}n_{1}^{\prime}n_{2}^{\prime},\ \ C=h_{1}^{*}n_{2}^{\prime}-h_{2}^{*}n_{1}^{\prime},\ \ ABC(AB-Cqt)\neq 0.$
Therefore, for any given $n_{1}^{\prime},n_{2}^{\prime},h_{1},h_{2},t$ therein and $h_{1}^{*},h_{2}^{*}$ as above, there exist at most $\tau(|AB-Cqt|)\ll(qHN)^{\varepsilon}$ choices of $n_{1},n_{2}.$ It then follows that
$\mathcal{B}_{2}$ $\ll H^{2}Tq^{\varepsilon}\mathop{\sum\sum}_{n_{1}^{\prime},n_{2}^{\prime}\sim N}(1+(n_{1}^{\prime},n_{2}^{\prime})H/N)$
$\ll(1+H/N)(1+HN^{3}/q)(HN)^{2}q^{\varepsilon}$
as desired. ∎
We now come back to the estimate for $\varSigma^{*}.$ From (6.8), (6.9) and Lemma 6.1, it follows that
$\varSigma^{*}$ $\ll\frac{M}{\sqrt{q}}(HN)^{2-\frac{1}{\nu}}(1+H/N)^{\frac{1}{2\nu}}(1+HN^{3}/q)^{\frac{1}{2\nu}}(K^{2\nu}q^{\frac{1}{2}}+K^{\nu}q)^{\frac{1}{2\nu}}q^{\varepsilon}.$
Recall that $H=x^{-1}MNq^{1+\varepsilon}$, $K=M^{-1}q^{1+\varepsilon}$ and $M\leqslant q^{-1}x^{1-\varepsilon}$. It is useful to make a restriction $N\ll q^{\frac{1}{4}}$, in which case we find $H\ll N$ and $HN^{3}\ll q^{1+\varepsilon}.$ Hence
$\varSigma^{*}$ $\ll(M^{2-\frac{1}{\nu}}q^{\frac{5}{2}-\frac{3}{4\nu}}+M^{\frac{5}{2}-\frac{1}{\nu}}q^{2-\frac{1}{2\nu}})x^{-2+\frac{1}{\nu}}N^{4-\frac{2}{\nu}}q^{\varepsilon}.$
Taking $\nu=3$, we arrive at
$\varSigma^{*}$ $\ll(M^{\frac{5}{3}}q^{\frac{9}{4}}+M^{\frac{13}{6}}q^{\frac{11}{6}})x^{-\frac{5}{3}}N^{\frac{10}{3}}q^{\varepsilon}.$
Together with (6.3) and (6.5), this yields
$\mathcal{R}(M,N)^{2}$ $\ll(xqM+x^{\frac{1}{3}}q^{\frac{9}{4}}M^{\frac{2}{3}}N^{\frac{4}{3}}+x^{\frac{1}{3}}q^{\frac{11}{6}}M^{\frac{7}{6}}N^{\frac{4}{3}}+qM^{2}N^{2})q^{\varepsilon}.$
Hence (6.1) holds for
(6.12) $M\leqslant q^{-1}x^{1-\varepsilon},\ \ N\leqslant\min\{x^{\frac{3}{4}}q^{-\frac{19}{16}},~{}x^{\frac{3}{8}}q^{-\frac{1}{2}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
since $q\geqslant x^{\frac{1}{2}}$.
Taking $\nu=4$, we arrive at
$\varSigma^{*}$ $\ll(M^{\frac{7}{4}}q^{\frac{37}{16}}+M^{\frac{9}{4}}q^{\frac{15}{8}})x^{-\frac{7}{4}}N^{\frac{7}{2}}q^{\varepsilon}.$
Together with (6.3) and (6.5), this yields
$\mathcal{R}(M,N)^{2}$ $\ll(xqM+x^{\frac{1}{4}}q^{\frac{37}{16}}M^{\frac{3}{4}}N^{\frac{3}{2}}+x^{\frac{1}{4}}q^{\frac{15}{8}}M^{\frac{5}{4}}N^{\frac{3}{2}}+qM^{2}N^{2})q^{\varepsilon}.$
Hence (6.1) holds for
(6.13) $M\leqslant q^{-1}x^{1-\varepsilon},\ \ N\leqslant\min\{x^{\frac{2}{3}}q^{-\frac{25}{24}},~{}x^{\frac{1}{3}}q^{-\frac{5}{12}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
since $q\geqslant x^{\frac{1}{2}}$.
6.4. Bounding $\varSigma^{*}$ as trilinear forms
We infer from (6.7) and Cauchy–Schwarz inequality that
(6.14) $\varSigma^{*}$ $\ll\frac{M}{\sqrt{q}}(\mathcal{T}_{1}\mathcal{T}_{2})^{\frac{1}{2}},$
where
$\mathcal{T}_{1}$ $=\sum_{b\leqslant B}\sum_{c\leqslant C}\rho(b,c)^{2},$
$\mathcal{T}_{2}$ $=\mathop{\sum\sum}_{\begin{subarray}{c}b,c\in\mathbf{Z}\\ (c,q)=1\end{subarray}}W\Big{(}\frac{b}{B}\Big{)}W\Big{(}\frac{c}{C}\Big{)}\Big{|}\sum_{1\leqslant|k|\leqslant K}\delta_{k}\mathrm{Kl}(akb\overline{c},q)\Big{|}^{2}.$
Squaring out and switching summations,
$\mathcal{T}_{2}$ $\ll\mathop{\sum\sum}_{1\leqslant|k_{1}|,|k_{2}|\leqslant K}|\mathcal{T}_{2}(k_{1},k_{2})|,$
where
$\mathcal{T}_{2}(k_{1},k_{2})$ $=\mathop{\sum\sum}_{\begin{subarray}{c}b,c\in\mathbf{Z}\\ (c,q)=1\end{subarray}}W\Big{(}\frac{b}{B}\Big{)}W\Big{(}\frac{c}{C}\Big{)}\mathrm{Kl}(ak_{1}b\overline{c},q)\mathrm{Kl}(ak_{2}b\overline{c},q).$
A trivial bound reads $\mathcal{T}_{2}(k_{1},k_{2})\ll BC$, which will be applied for diagonal terms. For off-diagonal terms, i.e., $k_{1}\neq k_{2}$, we would like to bound $\mathcal{T}_{2}(k_{1},k_{2})$ by appealing to the work of Fouvry, Kowalski and Michel [FKM14] on non-correlations of divisor functions and Frobenius trace functions; see Lemma 4.3.
From Poisson summation, it follows that
$\sum_{b\in\mathbf{Z}}W\Big{(}\frac{b}{B}\Big{)}\mathrm{Kl}(ak_{1}b\overline{c},q)\mathrm{Kl}(ak_{2}b\overline{c},q)$ $=\frac{B}{\sqrt{q}}\sum_{l\in\mathbf{Z}}\widehat{W}\Big{(}\frac{Bl}{q}\Big{)}K(lc,q;k_{1},k_{2}),$
where
$K(z,q;k_{1},k_{2})=\frac{1}{\sqrt{q}}\sum_{u\,(\operatorname{mod}{q})}\mathrm{Kl}(k_{1}u,q)\mathrm{Kl}(k_{2}u,q)\mathrm{e}\Big{(}\frac{-zu}{q}\Big{)}.$
This gives
$\mathcal{T}_{2}(k_{1},k_{2})$ $=\frac{B}{\sqrt{q}}\mathop{\sum\sum}_{\begin{subarray}{c}l,c\in\mathbf{Z}\\ (c,q)=1\end{subarray}}\widehat{W}\Big{(}\frac{Bl}{q}\Big{)}W\Big{(}\frac{c}{C}\Big{)}K(lc,q;k_{1},k_{2}).$
Observe that
$K(0,q;k_{1},k_{2})$ $=\frac{1}{\sqrt{q}}\sum_{u\,(\operatorname{mod}{q})}\mathrm{Kl}(k_{1}u,q)\mathrm{Kl}(k_{2}u,q)=q^{-\frac{1}{2}}S(k_{1}-k_{2};0;q).$
We employ the trivial bound for $l=0$, and appeal to Lemma 4.3 for $l\neq 0$ with the aid of Lemma 4.2. Hence the off-diagonal terms contribute at most
$\mathcal{T}_{2}(k_{1},k_{2})$ $\ll\frac{B}{q}(k_{1}-k_{2},q)+\frac{B}{\sqrt{q}}\frac{qC}{B}\Big{(}\frac{B}{C}+1\Big{)}^{\frac{1}{2}}q^{-\frac{1}{8}+\varepsilon}$
$\ll\frac{HN}{q}(k_{1}-k_{2},q)+\Big{(}\frac{H}{N}+1\Big{)}^{\frac{1}{2}}N^{2}q^{\frac{3}{8}+\varepsilon}.$
Summing over $k_{1},k_{2}$, we find
$\mathcal{T}_{2}$ $\ll HN^{3}K+\frac{HNK^{2}}{q}+\Big{(}\frac{H}{N}+1\Big{)}^{\frac{1}{2}}(NK)^{2}q^{\frac{3}{8}+\varepsilon}$
(6.15) $\ll\frac{q^{2+\varepsilon}N^{4}}{x}+\frac{(qN)^{2+\varepsilon}}{xM}+\frac{N^{2}}{M^{2}}q^{\frac{19}{8}+\varepsilon}$
since $M<x/q.$
Note that
$\mathcal{T}_{1}$ $\ll\mathop{\sum\sum\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}h_{1},h_{2},h_{1}^{\prime},h_{2}^{\prime}\leqslant H,~{}n_{1},n_{2},n_{1}^{\prime},n_{2}^{\prime}\sim N\\ h_{1}n_{2}-h_{2}n_{1}=h_{1}^{\prime}n_{2}^{\prime}-h_{2}^{\prime}n_{1}^{\prime},~{}n_{1}n_{2}=n_{1}^{\prime}n_{2}^{\prime}\end{subarray}}1\ll(HN)^{2}\Big{(}\frac{H}{N}+1\Big{)}q^{\varepsilon}\ll\Big{(}\frac{qMN^{2}}{x}\Big{)}^{2}q^{\varepsilon},$
from which and (6.15), (6.14) we derive that
$\varSigma^{*}$ $\ll\Big{(}\frac{q^{\frac{3}{2}}M^{2}N^{4}}{x^{\frac{3}{2}}}+\frac{(qM)^{\frac{3}{2}}N^{4}}{x^{\frac{3}{2}}}+\frac{q^{\frac{27}{16}}MN^{3}}{x}\Big{)}q^{\varepsilon}.$
Together with (6.3) and (6.5), this yields
$\mathcal{R}(M,N)^{2}$ $\ll(xqM+x^{\frac{1}{2}}q^{\frac{3}{2}}MN^{2}+x^{\frac{1}{2}}q^{\frac{3}{2}}M^{\frac{1}{2}}N^{2}+xq^{\frac{27}{16}}N+qM^{2}N^{2})q^{\varepsilon}.$
Hence (6.1) holds for
(6.16) $M\leqslant q^{-1}x^{1-\varepsilon},\ \ N\leqslant\min\{xq^{-\frac{27}{16}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon}$
since $q\geqslant x^{\frac{1}{2}}$.
In view of (6.12), (6.13) and (6.16), we find (6.1) is valid if taking $M=q^{-1}x^{1-\varepsilon}$ and
$N=\begin{cases}\min\{xq^{-\frac{27}{16}},~{}(x/q)^{\frac{1}{4}}\}\}x^{-\varepsilon},\ \ &\varpi\in[1/2,~{}32/61[,\\ \min\{x^{\frac{2}{3}}q^{-\frac{25}{24}},~{}x^{\frac{1}{3}}q^{-\frac{5}{12}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon},&\varpi\in[32/61,~{}7/13[,\\ \min\{x^{\frac{3}{4}}q^{-\frac{19}{16}},~{}x^{\frac{3}{8}}q^{-\frac{1}{2}},~{}(x/q)^{\frac{1}{4}}\}x^{-\varepsilon},&\varpi\in[7/13,~{}4/7[.\end{cases}$
Taking $D=MN$ in Proposition 5.1, we conclude the proof of Theorem 1.2.
Remark 3.
The estimates for diagonal terms become better when we take larger $\nu$ inHölder’s inequality. In particular, if we take $\nu=2$, the diagonal terms will lead to a constraint $MN\leqslant x^{\frac{3}{2}}q^{-\frac{7}{4}}$, which coincides with Iwaniec’s result. On the other hand, estimates for off-diagonal terms become worse when we take larger $\nu$.
6.5. Proof of Theorem 1.4
Recall that
$\mathcal{R}(M,N)^{2}$ $\ll x^{2}HN^{-1}q^{\varepsilon}+\frac{x^{2}}{MN^{2}}\varSigma$
with
$\varSigma$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\\ (n_{1}n_{2},q)=1\end{subarray}}\beta_{h_{1},n_{1},t}\overline{\beta_{h_{2},n_{2},t}}\sum_{\begin{subarray}{c}m\in\mathbf{Z}\\ (m,q)=1\end{subarray}}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}m}}{q}\Big{)}.$
as in (6.3) and (6.4). Applying Lemma 4.7 with $h=a(h_{1}n_{2}-h_{2}n_{1})\overline{n_{1}n_{2}}$ and partial summation to the inner sum, we get
$\varSigma$ $\ll\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\end{subarray}}\Big{(}\frac{M}{q}(h_{1}n_{2}-h_{2}n_{1},q)+q^{\kappa}M^{\lambda-\kappa+{O(\eta)}}(h_{1}n_{2}-h_{2}n_{1},q)^{\nu}\Big{)}.$
Note that
$\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}0<|h_{1}|,|h_{2}|\leqslant H,~{}n_{1},n_{2}\sim N\\ h_{1}n_{2}\neq h_{2}n_{1}\end{subarray}}(h_{1}n_{2}-h_{2}n_{1},q)$ $\ll(HN)^{2}q^{\varepsilon},$
we thus obtain
$\varSigma$ $\ll(x^{-2}M^{3}N^{4}q+x^{-2}M^{2+\lambda-\kappa}N^{4}q^{\kappa+2})q^{O(\eta)},$
which yields
$\mathcal{R}(M,N)\leqslant(x^{\frac{1}{2}}q^{\frac{1}{2}}M^{\frac{1}{2}}+q^{\frac{1}{2}}MN+q^{1+\frac{\kappa}{2}}M^{\frac{1-\kappa+\lambda}{2}}N)q^{O(\eta)}.$
Hence (6.1) holds for
$M=q^{-1}x^{1-\varepsilon},\ \ N=\min\{x^{\frac{1+\kappa-\lambda}{2}}q^{-\frac{1+2\kappa-\lambda}{2}},~{}q^{\frac{1}{2}}\}x^{-\varepsilon}$
with $\varepsilon\gg\eta.$ In particular, for $\varpi\geqslant\frac{1+\kappa-\lambda}{2+2\kappa-\lambda}$, we may take
$D=x^{\frac{3+\kappa-\lambda}{2}-2\varepsilon}q^{-\frac{3+2\kappa-\lambda}{2}}$
in Proposition 5.1. Note that van Lint and Richert [LR65] obtained a level $q^{-1}x^{1-\varepsilon}$, and the above choice of $D$ is better as long as $\varpi\leqslant\frac{1+\kappa-\lambda}{1+2\kappa-\lambda}$. The proof of Theorem 1.4 is now finished.
7. Proof of Theorem 3.1
7.1. Preparations
Denote by $\omega(n)$ the number of distinct prime factors of $n$. We first recall a very crucial observation due to Fouvry [Fo84, Lemme 6].
Lemma 7.1.
For $\xi\in\mathbf{Z}^{+}$ and $M,N\geqslant 2,$ denote
$\mathscr{F}_{\xi}(M,N)$ =$=\{(m,n)\in[1,M]\times[1,N]:(m,n)=1\text{ and }\omega(m),\omega(n)\leqslant\xi\}.$
Then there exists a partition of $\mathscr{F}_{\xi}(M,N)$ into at most $(\log(MN))^{2\xi}$ subsets $\mathscr{F}^{*}_{\xi}(M,N)$ having the following property:
$(m_{1},n_{1}),(m_{2},n_{2})\in\mathscr{F}^{*}_{\xi}(M,N)\Longrightarrow(m_{1},n_{2})=1.$
[…]
(7.1) $-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{(mn,rs)=(ns,a)=1}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)}.$
The proof of Theorem 3.1 would require the following estimate.
Lemma 7.2.
[…]
$-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $\ll\|\boldsymbol{\alpha}\|(HNS)^{\frac{1}{2}}((HNS)^{\frac{1}{4}}+(NS)^{\frac{1}{2}})(\log(NS))^{2\xi}\Bigg{[}NS(MR)^{\frac{1}{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}$
×$\ \ \ \ \times\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}$
$\ \ \ \ +\|\boldsymbol{\alpha}\|(HMNRS)^{\frac{1}{2}+\varepsilon}(\log(NS))^{2\xi},$
where the implied constant depends only on $\varepsilon,$ and $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
Proof.
According to Lemma 7.1, we split the sums over $n,s$ into at most $O((\log NS)^{2\xi})$ sub-sums of the shape
$\mathcal{Q}^{*}=\mathcal{Q}_{a}^{*}\{\mathscr{F}^{*}_{\xi}(N,S)\}$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{\begin{subarray}{c}(mn,rs)=(ns,a)=1\\ (n,s)\in\mathscr{F}^{*}_{\xi}(N,S)\end{subarray}}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)},$
so that
(7.2) $-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{\mathscr{F}^{*}_{\xi}(N,S)}\mathcal{Q}_{a}^{*}\{\mathscr{F}^{*}_{\xi}(N,S)\}.$
By Cauchy–Schwarz inequality, we have
$|\mathcal{Q}^{*}|^{2}$ $\leqslant\|\boldsymbol{\alpha}\|^{2}\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (r,m)=1\end{subarray}}W\Big{(}\frac{r}{R}\Big{)}W\Big{(}\frac{m}{M}\Big{)}\Big{|}\mathop{\sum_{h\leqslant H}\sum_{n\leqslant N}\sum_{s\leqslant S}}_{\begin{subarray}{c}(n,rs)=(ns,a)=1\\ (n,s)\in\mathscr{F}^{*}_{\xi}(N,S)\end{subarray}}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)}\Big{|}^{2}.$
Squaring out and switching summations,
$|\mathcal{Q}^{*}|^{2}$ $\leqslant\|\boldsymbol{\alpha}\|^{2}\mathop{\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}h_{1},h_{2}\leqslant H,~{}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1,~{}(n_{1},s_{1}),(n_{2},s_{2})\in\mathscr{F}^{*}_{\xi}(N,S)\end{subarray}}\beta_{h_{1},n_{1},s_{1}}\overline{\beta_{h_{2},n_{2},s_{2}}}$
×$\ \ \ \ \times\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (ar,n_{1}n_{2})=(am,s_{1}s_{2})=(r,m)=1\end{subarray}}W\Big{(}\frac{r}{R}\Big{)}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{ah_{1}\overline{mn_{1}}}{rs_{1}}-\frac{ah_{2}\overline{mn_{2}}}{rs_{2}}\Big{)}.$
The diagonal terms, i.e., $h_{1}n_{2}s_{2}=h_{2}n_{1}s_{1},$ give a contribution at most
$\ll\|\boldsymbol{\alpha}\|^{2}MR\mathop{\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}h_{1},h_{2}\leqslant H,~{}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ h_{1}n_{2}s_{2}=h_{2}n_{1}s_{1}\end{subarray}}1$
$\ll\|\boldsymbol{\alpha}\|^{2}HMNRS(HNS)^{\varepsilon}.$
Denote by $\mathcal{Q}^{*}$ the contribution from off-diagonal terms, so that
(7.3) $|\mathcal{Q}^{*}|^{2}$ $\leqslant\mathcal{Q}_{1}+O(\|\boldsymbol{\alpha}\|^{2}HMNRS(HNS)^{\varepsilon}).$
For tuples $(n_{1},s_{1}),(n_{2},s_{2})\in\mathscr{F}_{\xi}^{*}(N,S),$ we find $n_{1},n_{2},s_{1},s_{2}$ are pairwise coprime according to Lemma 7.1. Hence the above fraction in the exponential is
$\frac{a\overline{mn_{1}n_{2}}(h_{1}n_{2}s_{2}-h_{2}n_{1}s_{1})}{rs_{1}s_{2}}\,(\operatorname{mod}{1}).$
For given $n_{1},n_{2},s_{1},s_{2}$, we introduce a coefficient $\boldsymbol{\gamma}_{n_{1},n_{2},s_{1},s_{2}}=(\gamma_{k;n_{1},n_{2},s_{1},s_{2}})$, indexed by $k$, as
$\gamma_{k;n_{1},n_{2},s_{1},s_{2}}$ $=\mathop{\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}h_{1},h_{2}\leqslant H,~{}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1,~{}(n_{1},s_{1}),(n_{2},s_{2})\in\mathscr{F}_{\xi}^{*}(N,S)\\ h_{1}n_{2}s_{2}-h_{2}n_{1}s_{1}=k\end{subarray}}\beta_{h_{1},n_{1},s_{1}}\overline{\beta_{h_{2},n_{2},s_{2}}}.$
This convention allows us to write
$\mathcal{Q}_{1}$ $=\|\boldsymbol{\alpha}\|^{2}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1,~{}(n_{1},s_{1}),(n_{2},s_{2})\in\mathscr{F}_{\xi}^{*}(N,S)\end{subarray}}\sum_{1\leqslant|k|\leqslant HNS}\gamma_{k;n_{1},n_{2},s_{1},s_{2}}$
×$\ \ \ \ \times\mathop{\sum\sum}_{\begin{subarray}{c}r,m\in\mathbf{Z}\\ (ar,n_{1}n_{2})=(am,s_{1}s_{2})=(r,m)=1\end{subarray}}W\Big{(}\frac{r}{R}\Big{)}W\Big{(}\frac{m}{M}\Big{)}\mathrm{e}\Big{(}\frac{ak\overline{mn_{1}n_{2}}}{rs_{1}s_{2}}\Big{)}.$
We are now in a position to apply Lemma 3.3 with
$(q,r,s)\leftarrow(a,n_{1}n_{2},s_{1}s_{2}),\ \ (c,m,n)\leftarrow(r,m,k),\ \ (C,M,N)\leftarrow(R,M,HNS),$
getting
(7.4) $\mathcal{Q}_{1}$ $\ll\|\boldsymbol{\alpha}\|^{2}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1\end{subarray}}\|\boldsymbol{\gamma}_{n_{1},n_{2},s_{1},s_{2}}\|\cdot\Delta,$
where
$\Delta$ $=|a|^{\vartheta}\sqrt{MRS^{2}N^{2}}\frac{(1+\frac{MRS^{2}N^{2}}{|a|HNS})^{\vartheta}}{1+(\frac{|a|HNS}{MRS^{2}N^{2}})^{2}}\Big{(}1+\frac{|a|HNS}{MRS^{2}N^{2}}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}$
×$\ \ \ \times\Big{(}1+\frac{|a|HNS}{MRS^{2}N^{2}}+\frac{HNS}{N^{2}S^{2}}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}$
$=\sqrt{MN^{2}RS^{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}.$
Lemma 7.3.
$\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1\end{subarray}}\|\boldsymbol{\gamma}_{n_{1},n_{2},s_{1},s_{2}}\|\ll(H(NS)^{2}+(HNS)^{\frac{3}{2}})(HNS)^{\varepsilon}.$
Proof.
We first observe that
$\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (n_{1}n_{2},s_{1}s_{2})=1\end{subarray}}\|\boldsymbol{\gamma}_{n_{1},n_{2},s_{1},s_{2}}\|^{2}$ $\ll\mathop{\sum\sum\sum\sum\sum\sum\sum\sum}_{\begin{subarray}{c}h_{1},h_{2},h_{1}^{\prime},h_{2}^{\prime}\leqslant H,~{}n_{1},n_{2}\leqslant N,~{}s_{1},s_{2}\leqslant S\\ (h_{1}-h_{1}^{\prime})n_{2}s_{2}=(h_{2}-h_{2}^{\prime})n_{1}s_{1}\end{subarray}}1$
$\ll((HNS)^{2}+H^{3}NS)(HNS)^{\varepsilon}.$
The lemma then follows from the Cauchy–Schwarz inequality. ∎
From (7.4) and Lemma 7.3, it follows that
$\mathcal{Q}_{1}$ $\ll\|\boldsymbol{\alpha}\|^{2}HNS((HNS)^{\frac{1}{2}}+NS)\Bigg{[}NS(MR)^{\frac{1}{2}}\frac{(|a|+\frac{MNRS}{H})^{\vartheta}}{1+(\frac{|a|H}{MNRS})^{2}}\Big{(}1+\frac{|a|H}{MNRS}+\frac{R}{MN^{2}}\Big{)}^{\frac{1}{2}}$
×$\ \ \ \ \times\Big{(}1+\frac{|a|H}{MNRS}+\frac{H}{NS}\Big{)}^{\frac{1}{2}}+S^{-2}M(HNS)^{\frac{1}{2}}\Bigg{]},$
from which and (7.3), (7.2) we conclude Lemma 7.2. ∎
7.2. Initial reductions
[…]
$2^{\xi}\gg(\log(MNq))^{A},\ \ \ (\log N)^{\xi}\ll\exp(\xi^{2})\ll(MNq)^{\varepsilon}$
for any $A,\varepsilon>0.$
We first consider a sub-sum
$\mathcal{R}_{0}:=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M,~{}n_{1},n_{2}\in\mathbf{Z}\\ m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})\\ (m_{1}m_{2}n_{1}n_{2},q)=1,~{}\omega(m_{1})>\xi\end{subarray}}\alpha_{m_{1}}\beta_{m_{2}}F(n_{1},n_{2}).$
Trivially, we have
$\mathcal{R}_{0}$ $\ll\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M,~{}n_{1},n_{2}\sim N\\ m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})\\ \omega(m_{1})>\xi\end{subarray}}1\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}m_{1}\sim M,~{}n_{1}\sim N\\ (m_{1}n_{1},q)=1,~{}\omega(m_{1})>\xi\end{subarray}}\sum_{\begin{subarray}{c}MN<k\leqslant 4MN\\ k\equiv m_{1}n_{1}\,(\operatorname{mod}{q})\end{subarray}}\tau(k).$
By Shiu’s bound (see [Sh80, Theorem 2]) we obtain
$\sum_{k}\ll\frac{MN\log(MN)}{\varphi(q)}.$
Moreover,
$\sum_{\begin{subarray}{c}m_{1}\sim M\\ \omega(m_{1})>\xi\end{subarray}}1$ $\leqslant\frac{1}{2^{\xi}}\sum_{m_{1}\sim M}2^{\omega(m)}\ll\frac{M\log M}{2^{\xi}}\ll M(\log MNq)^{-A-2}$
for any $A>0$. Hence
$\mathcal{R}_{0}\ll q^{-1}(MN)^{2}(\log MNq)^{-A}.$
In this way, we may write
(7.5) =~$\mathcal{R}=\widetilde{\mathcal{R}}+O(q^{-1}(MN)^{2}(\log MNq)^{-A}),$
where $\widetilde{\mathcal{R}}$ is defined as $\mathcal{R},$ with an additional restriction that
$\omega(m_{1})>\xi\text{ or }\omega(m_{2})>\xi\Longrightarrow\alpha_{m_{1}}\beta_{m_{2}}=0,$
i.e., we additionally assume that $(m_{1},m_{2})\in\mathscr{F}_{\xi}(2M,2M).$
We now turn to estimate $\widetilde{\mathcal{R}}.$ We first write
~$\widetilde{\mathcal{R}}=\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d,~{}n_{1},n_{2}\in\mathbf{Z}\\ m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1},m_{2})=1\end{subarray}}\alpha_{dm_{1}}\beta_{dm_{2}}F(n_{1},n_{2}),$
whence we assume that $(dm_{1},dm_{2})\in\mathscr{F}_{\xi}(2M,2M).$ Note that the congruence restriction $m_{1}n_{1}\equiv m_{2}n_{2}\,(\operatorname{mod}{q})$ is equivalent to
[…]
with $R:=2MN/q.$ Therefore,
$\widetilde{\mathcal{R}}$ $=\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\alpha_{dm_{1}}\beta_{dm_{2}}$
×$\ \ \ \ \times\sum_{\begin{subarray}{c}n_{1}\equiv qr\overline{m_{1}}\,(\operatorname{mod}{m_{2}})\\ (n_{1},q)=1\end{subarray}}F\Big{(}n_{1},\frac{m_{1}n_{1}-qr}{m_{2}}\Big{)}.$
From Poisson summation, the last $n_{1}$-sum can be transformed as
$\sum_{n_{1}}$ $=\sum_{\delta\mid q}\mu(\delta)\sum_{n_{1}\equiv qr\overline{\delta m_{1}}\,(\operatorname{mod}{m_{2}})}F\Big{(}\delta n_{1},\frac{\delta m_{1}n_{1}-qr}{m_{2}}\Big{)}$
$=\frac{1}{m_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta m_{1}}}{m_{2}}\Big{)},$
where
$\widetilde{F}(\lambda;m_{1},m_{2},r)=\int_{\mathbf{R}}F\Big{(}y,\frac{m_{1}y-qr}{m_{2}}\Big{)}\mathrm{e}(-\lambda y)\mathrm{d}y.$
Note that the integration by parts yields
(7.6) $\sup_{m_{1},m_{2},r}|\widetilde{F}(\lambda;m_{1},m_{2},r)|$ $\ll N(1+\eta^{-1}|\lambda|N)^{-A}$
for all $A\geqslant 0$. Hence
$\widetilde{\mathcal{R}}$ $=\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{h\in\mathbf{Z}}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta m_{1}}}{m_{2}}\Big{)}.$
The term with $h=0$ gives exactly
$=\frac{\varphi(q)}{q}\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}\sum_{0\leqslant|r|\leqslant R/d}\widetilde{F}(0;m_{1},m_{2},r).$
Note that the $r$-sum is
$=m_{2}\sum_{r\in\mathbf{Z}}\int_{\mathbf{R}}F\Big{(}m_{2}y,m_{1}y-\frac{qr}{m_{2}}\Big{)}\mathrm{d}y$
$=m_{2}\iint_{\mathbf{R}^{2}}F\Big{(}m_{2}y,m_{1}y-\frac{qz}{m_{2}}\Big{)}\mathrm{d}y\mathrm{d}z+O(\eta q^{-100})$
$=\frac{m_{2}}{q}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z+O(\eta q^{-100}).$
Hence
(7.7) ~$\widetilde{\mathcal{R}}=\mathcal{R}^{*}+\frac{\varphi(q)}{q^{2}}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M\\ (m_{1}m_{2},q)=1\end{subarray}}\alpha_{m_{1}}\beta_{m_{2}}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z+O(\eta),$
where
$\mathcal{R}^{*}$ $=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{0\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}$
×$\ \ \ \ \times\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{qrh\overline{\delta m_{1}}}{m_{2}}\Big{)}.$
Denote by $\mathcal{R}^{*}_{1}$ the contribution to $\mathcal{R}^{*}$ from the term with $r=0$. Note that
$\sum_{h\neq 0}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},0\Big{)}$ $\ll\eta N^{-1+\varepsilon}\delta m_{2}$
by (7.6). Hence
$\mathcal{R}^{*}_{1}\ll\eta(MN)^{1+\varepsilon},$
so that
(7.8) $\mathcal{R}^{*}$ $=\mathcal{R}^{*}_{2}+O(\eta(MN)^{1+\varepsilon}),$
where
$\mathcal{R}^{*}_{2}$ $=\sum_{\delta\mid q}\frac{\mu(\delta)}{\delta}\sum_{\begin{subarray}{c}d\leqslant 2M\\ (d,q)=1\end{subarray}}\sum_{1\leqslant|r|\leqslant R/d}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M/d\\ (m_{1}m_{2},q)=(m_{1},m_{2})=1\end{subarray}}\frac{\alpha_{dm_{1}}\beta_{dm_{2}}}{m_{2}}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H\delta/d}\widetilde{F}\Big{(}\frac{h}{m_{2}\delta};m_{1},m_{2},r\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}}}{m_{2}}\Big{)}$
with $H:=\eta M^{1+\varepsilon}/N.$
7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
[…]
$\mathcal{R}^{*}_{2}$ $\ll(MNq)^{\varepsilon}\sum_{\delta\mid q}\frac{1}{\delta}\sum_{\begin{subarray}{c}d_{1}d_{2}=d\leqslant 2M\\ (d,q)=1\end{subarray}}\sup|\mathcal{R}^{*}_{2}(A_{1},B_{1},A_{2},B_{2},R^{\prime};d_{1},d_{2},\delta)|,$
where the supremum is over
$1\leqslant A_{1}\leqslant\frac{M_{1}}{d_{1}},~{}1\leqslant B_{1}\leqslant\frac{N_{1}}{d_{2}},~{}1\leqslant A_{2}\leqslant\frac{M_{2}}{d_{1}},~{}1\leqslant B_{2}\leqslant\frac{N_{2}}{d_{2}},~{}1\leqslant R^{\prime}\leqslant\frac{R}{d},$
and
$\mathcal{R}^{*}_{2}(\cdots)$ $=\sum_{|r|\sim R^{\prime}}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=1\end{subarray}}\frac{\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}m_{2}}\beta^{\prime\prime}_{d_{2}n_{2}}}{m_{2}n_{2}}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H\delta/d}\widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big{)}.$
We would like explore the cancellations among the exponentials $\mathrm{e}(\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}})$ using Lemma 7.2. It is now necessary to separate variables. By Fourier inversion,
$\ \ \ \ \widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}$
$=m_{2}n_{2}\delta\int_{\frac{dN}{2M\delta}}^{\frac{2dN}{M\delta}}F\Big{(}m_{2}n_{2}y\delta,m_{1}n_{1}y\delta-\frac{qr}{m_{2}n_{2}}\Big{)}\mathrm{e}(-hy)\mathrm{d}y$
$=m_{2}n_{2}\delta\int_{\frac{dN}{2M\delta}}^{\frac{2dN}{M\delta}}\mathrm{e}(-hy)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}(y_{1},y_{2})\mathrm{e}\Big{(}m_{2}n_{2}y_{1}y\delta+m_{1}n_{1}y_{2}y\delta-\frac{qry_{2}}{m_{2}n_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
$=\frac{m_{2}n_{2}\delta}{r^{2}}\int_{\frac{dN}{2M\delta}}^{\frac{2dN}{M\delta}}\mathrm{e}(-hy)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}\Big{(}\frac{m_{1}y_{1}}{n_{2}r},\frac{y_{2}n_{2}}{m_{1}r}\Big{)}\mathrm{e}\Big{(}\frac{(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta}{r}-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
$=\frac{m_{2}n_{2}\delta}{r}\int_{\frac{dN}{2Mr\delta}}^{\frac{2dN}{Mr\delta}}\mathrm{e}(-hry)\mathrm{d}y\iint_{\mathbf{R}^{2}}\widehat{F}\Big{(}\frac{m_{1}y_{1}}{n_{2}r},\frac{y_{2}n_{2}}{m_{1}r}\Big{)}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}.$
Note that for all $\nu_{1},\nu_{2}\in\mathbf{N},$
(7.9) $\frac{\partial^{\nu_{1}+\nu_{2}}}{\partial y_{1}^{\nu_{1}}\partial y_{2}^{\nu_{2}}}\widehat{F}(y_{1},y_{2})\ll N^{\nu_{1}+\nu_{2}+2}(1+\eta^{-1}N|y_{1}|)^{-c_{1}}(1+\eta^{-1}N|y_{2}|)^{-c_{2}}$
with any $c_{1},c_{2}\geqslant 0$, and with an implied constant depending only on $(\nu_{1},\nu_{2},c_{1},c_{2})$.
We need further to separate variables in $\widehat{F}(*,*)$ by Mellin inversion. To this end, we truncate the integral over $y_{1}$ over five segments
$I_{1}=]-\infty,-V_{1}],~{}I_{2}=[-V_{1},-U],~{}I_{3}=[-U,U],~{}I_{4}=[U,V_{1}],~{}I_{5}=[V_{1},+\infty[$
with
$U=\frac{1}{M^{2}N^{4}},\ \ V_{1}=\frac{\eta R^{\prime}B_{2}}{A_{1}N}N^{\varepsilon}.$
In view of the decay (7.9) of $\widehat{F},$ we find the contributions from $I_{1},I_{5}$ are negligible, and we may also bound the integral over $I_{3}$ trivially, contributing to $\mathcal{R}^{*}_{2}(\cdots)$ at most
$\ll\frac{\eta\delta HA_{1}^{2}A_{2}B_{1}}{M^{3}N^{2}}\ll\frac{\eta^{2}\delta M^{\varepsilon}}{M^{2}N},$
which is also negligible. It thus suffices to treat the integral over $y_{1}\in I_{2}\cup I_{4}$. In a similar manner, we may also reduce the integral over $y_{2}$ to $[-V_{2},-U]\cup[U,V_{2}]$, where
$V_{2}=\frac{\eta R^{\prime}A_{1}}{B_{2}N}N^{\varepsilon}.$
By symmetry, we only transform the integrals over $y_{1}\in[U,V_{1}]$ and $y_{2}\in[U,V_{2}],$ and the corresponding contributions are still denoted by $\widetilde{F}(\cdots)$ for convenience.
Note that, for $w_{1},w_{2}>0,$ the Mellin inversion gives
$\widehat{F}(w_{1},w_{2})$ $=\frac{1}{4\pi^{2}}\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})w_{1}^{-it_{1}}w_{2}^{-it_{2}}\mathrm{d}t_{1}\mathrm{d}t_{2},$
where
$\widetilde{F}(s_{1},s_{2})$ $=\iint_{(\mathbf{R}^{+})^{2}}\widehat{F}(w_{1},w_{2})w_{1}^{s_{1}-1}w_{2}^{s_{2}-1}\mathrm{d}w_{1}\mathrm{d}w_{2}.$
Hence
$\ \ \ \ \widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}$
$=\frac{m_{2}n_{2}\delta}{4\pi^{2}r}\int_{\frac{dN}{4MR^{\prime}\delta}}^{\frac{4dN}{MR^{\prime}\delta}}\mathrm{e}(-hry)\mathrm{d}y\int_{U}^{V_{1}}\int_{U}^{V_{2}}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{d}y_{1}\mathrm{d}y_{2}$
×$\ \ \ \ \times\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})\Big{(}\frac{n_{2}r}{m_{1}y_{1}}\Big{)}^{it_{1}}\Big{(}\frac{m_{1}r}{y_{2}n_{2}}\Big{)}^{it_{2}}\mathrm{d}t_{1}\mathrm{d}t_{2}.$
Thanks to the separation of variables as above, we get
$\mathcal{R}^{*}_{2}(\cdots)$ $=\frac{\delta}{4\pi^{2}}\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})\mathrm{d}t_{1}\mathrm{d}t_{2}\int_{\frac{dN}{4MR^{\prime}\delta}}^{\frac{4dN}{MR^{\prime}\delta}}\mathrm{d}y\int_{U}^{V_{1}}y_{1}^{-it_{1}}\mathrm{d}y_{1}\int_{U}^{V_{2}}y_{2}^{-it_{2}}\mathrm{d}y_{2}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H\delta/d}\sum_{|r|\sim R^{\prime}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=1\end{subarray}}\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}m_{2}}\beta^{\prime\prime}_{d_{2}n_{2}}$
×$\ \ \ \ \times\Big{(}\frac{n_{2}}{m_{1}}\Big{)}^{i(t_{1}-t_{2})}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big{)}.$
Note that the integration by parts and (7.9) yield
$\widetilde{F}(it_{1},it_{2})\ll(N\log N)^{2}(1+\eta^{-1}|t_{1}|)^{-c_{1}}(1+\eta^{-1}|t_{2}|)^{-c_{2}}$
with any $c_{1},c_{2}\geqslant 0$. This implies
$\iint_{\mathbf{R}^{2}}|\widetilde{F}(it_{1},it_{2})|\mathrm{d}t_{1}\mathrm{d}t_{2}\ll(\eta N\log N)^{2}.$
We are now in a good position to apply Lemma 7.2 with
$(a,h,m,n,r,s)\leftarrow(q/\delta,k,m_{1},n_{1},m_{2},n_{2}),\ \ \ (H,M,N,R,S)\leftarrow(R^{\prime}H\delta/d,A_{1},B_{1},A_{2},B_{2}),$
$\alpha_{m,r}\leftarrow\alpha^{\prime}_{d_{1}m_{1}}\beta^{\prime}_{d_{1}m_{2}}m_{1}^{i(t_{2}-t_{1})}\mathrm{e}\Big{(}m_{1}m_{2}y_{1}y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)},$
$\beta_{h,n,s}\leftarrow\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime\prime}_{d_{2}n_{2}}n_{2}^{i(t_{1}-t_{2})}\mathrm{e}(n_{1}n_{2}y_{2}y\delta)\mathop{\sum\sum}_{\begin{subarray}{c}hr=k\\ 1\leqslant|h|\leqslant H\delta/d,~{}|r|\sim R^{\prime}\end{subarray}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1},$
so that
$\mathcal{R}^{*}_{2}$ $\ll(MNq)^{\varepsilon}\frac{\eta^{4}N}{M}(RHM^{2})^{\frac{1}{2}}((RHB_{1}B_{2})^{\frac{1}{4}}+(B_{1}B_{2})^{\frac{1}{2}})$
×$\ \ \ \ \times\Bigg{[}q^{\vartheta}B_{1}B_{2}(A_{1}A_{2})^{\frac{1}{2}}\Big{(}1+\frac{qRH}{M^{2}}+\frac{A_{2}}{MB_{1}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{qRH}{M^{2}}+\frac{RH}{B_{1}B_{2}}\Big{)}^{\frac{1}{2}}+B_{2}^{-2}A_{1}(RHB_{1}B_{2})^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}$
$\ \ \ \ +\eta^{4}NM^{-1}(A_{1}A_{2}RHM^{2})^{\frac{1}{2}+\varepsilon}$
$\ll(MNq)^{\varepsilon}\eta^{\frac{5}{2}}MNq^{-\frac{1}{2}}\Big{(}\Big{(}\frac{M^{2}B_{1}B_{2}}{q}\Big{)}^{\frac{1}{4}}+(B_{1}B_{2})^{\frac{1}{2}}\Big{)}$
×$\ \ \ \ \times\Bigg{[}q^{\vartheta}B_{1}B_{2}(A_{1}A_{2})^{\frac{1}{2}}\Big{(}1+\frac{M^{2}}{B_{1}B_{2}q}\Big{)}^{\frac{1}{2}}+B_{2}^{-2}A_{1}\Big{(}\frac{M^{2}B_{1}B_{2}}{q}\Big{)}^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}$
$\ \ \ \ +(MNq)^{\varepsilon}\eta^{\frac{5}{2}}MN(A_{1}A_{2}/q)^{\frac{1}{2}},$
Taking $B_{1}B_{2}=(1+q^{-1}M^{2})(MNq)^{4\varepsilon}$, we get
$\mathcal{R}^{*}_{2}$ $\ll(MNq)^{\varepsilon}\eta^{\frac{9}{2}}MNq^{-\frac{1}{2}}(1+Mq^{-\frac{1}{2}})[q^{\vartheta}B_{1}B_{2}(A_{1}A_{2})^{\frac{1}{2}}]^{\frac{1}{2}}$
$\ \ \ \ +(MNq)^{\varepsilon}\eta^{\frac{5}{2}}MN(A_{1}A_{2}/q)^{\frac{1}{2}}$
$\ll\eta^{\frac{9}{2}}M^{\frac{3}{2}}Nq^{\frac{\vartheta-1}{2}}(1+Mq^{-\frac{1}{2}})^{\frac{3}{2}}(MNq)^{3\varepsilon}+\eta^{\frac{9}{2}}MN(MNq)^{-\varepsilon},$
from which and (7.8), (7.7), (7.5), we then conclude the proof of Theorem 3.1 by noticing the arbitrariness of $\varepsilon.$
8. Comments and remarks
8.1. Brun–Titchmarsh theorem on average
Hooley [Ho72] was the first who was able to study upper bounds of $\pi(x;q,a)$ non-tivially through extra averaging over $q$. This was largely motivated by the study of greatest prime factors of shifted primes, as approximations to the twin prime conjecture. For an individual $q$, one may also consider the average over the residue class $a\,(\operatorname{mod}{q})$. See also [Mo74, Ho75, Mo79, Iw82, Fo85] for these developments.
8.2. A variant of Theorem 3.1
As one may see from the proof of Theorem 3.1, one needs to estimate the exponential sum
$\sum_{k\leqslant K}\delta_{k}\sum_{m\leqslant M}\sum_{n\leqslant N}\alpha_{m}\beta_{n}\mathrm{e}\Big{(}ak\frac{\overline{m}}{n}\Big{)}.$
Since the above $\alpha_{m},\beta_{n}$ both factorize into convolutions in the prerequisites of Theorem 3.1, it is in fact reasonable to bound the quintilinear form (7.1). The flexibilities in convolutions are quite crucial in the proof of Theorem 3.1, and this structure can be guaranteed by Lemma 2.1 as developed by Iwaniec.
For arbitrary coefficients $-β\boldsymbol{\alpha},\bbeta,$ Bettin and Chandee [BC18] obtained an estimate for the above trilinear form, as a successor of a bound due to Duke, Friedlander and Iwaniec [DFI97] with $K=1$.
Lemma 8.1.
[…]
×$\qquad\times((M+N)^{\frac{1}{4}}+(KMN)^{\frac{1}{40}}(KM+KN)^{\frac{1}{8}}).$
The arguments in Section 7, together with Lemma 8.1, yield the following alternative formulation of (3.2) for arbitrary coefficients.
Theorem 8.1.
[…]
where the implied constant depends only on $\varepsilon$.
We also recall that Friedlander and Iwaniec [FI92] studied the twisted second moment
$\mathcal{Z}:=\frac{1}{\varphi(q)}\sum_{\chi\,(\operatorname{mod}{q})}\Big{|}\sum_{m\leqslant M}c_{m}\chi(m)\Big{|}^{2}\Big{|}\sum_{l\leqslant L}\chi(l)\Big{|}^{2}$
with $|c_{m}|\ll m^{\varepsilon},$ which is intimately related to the counting function (3.1). We now compare [FI92], Theorems 3.1 and 8.1, with an expectation that an asymptotic formula holds for $\mathcal{Z}$ with $M$ as large as possible. Take $L=q^{1/2}$ for interest.
  • Friedlander and Iwaniec [FI92] succeeded for $M\leqslant q^{2/7-}$ and $(c_{m})$ is a convolution of two coefficients of level $q^{1/7-}.$
  • Theorem 3.1 succeeds for $M\leqslant q^{(3-2\vartheta)/4-}$ and $c_{m}$ is well-factorable.
  • Theorem 8.1 succeeds for $M\leqslant q^{2/3-}$ and $c_{m}$ is general.
Recall the quintilinear form in (7.1):
$-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{(mn,rs)=(ns,a)=1}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)},$
which appears in the proofs of Theorem 3.1 and [FI92, Theorem 1]. In particular, due to the application of Kloostermania, the size of $a$ plays an important role in the settings of Theorem 3.1. Note that Lemma 8.1 follows from a careful application of amplification, so that only elementary means together with Weil’s bound for complete algebraic exponential sums are needed. A natural problem is to ask for an elementary bound for $-β\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ which is uniform in $a\neq 0$, and $-β\boldsymbol{\alpha},\bbeta$ could be well-factorable if necessary.
8.3. Twisted moments of Dirichlet $L$-functions
As one may see from the above arguments, it is important to study the twisted moment
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m\leqslant M}c_{m}\chi(m)\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{2k},$
and prove an averaged version of Lindelöf Hypothesis with $M$ as large as possible. We look forward to more advanced techniques to enter the picture. For $t=0,k=2$ and $M=q^{1/4}$ with $q$ prime, we refer to the work of Blomer, Humphries, Khan and Milinovich [BHKM20]. This is very cogent to make us believe that spectral reciprocity contributes as a very key input, so that the resultant bound is independent from the Ramanujan–Petersson conjecture. As mentioned before, it is highly desired to extend their bound to arbitrary moduli $q$.
8.4. Hooley’s Hypothesis R and impacts
In his studies on the Brun–Titchmarsh theorem and greatest prime factors of cubic polynomials, Hooley [Ho72, Ho78] proposed the following conjecture:
Conjecture 8.1 (Hypothesis R).
Let $q$ be a positive integer and $h\in\mathbf{Z}.$ Let $I$ be an interval with $1<|I|\leqslant q.$ Then
$\sum_{\begin{subarray}{c}n\in I\\ (n,q)=1\end{subarray}}\mathrm{e}\Big{(}\frac{h\overline{n}}{q}\Big{)}$ $\ll|I|^{\frac{1}{2}}(h,q)^{\frac{1}{2}}q^{\varepsilon}$
for any $\varepsilon>0.$
Under Hypothesis R, Iwaniec [Iw82, Theorems 8 & 9] showed the choices
(8.1) $C(\varpi)=\begin{cases}6/(5-6\varpi),\quad&\varpi\in]4/9,~{}7/12[,\\ 5/(3-3\varpi),&\varpi\in]7/12,~{}1[,\end{cases}$
are admissible in (1.3).
Hypothesis R exhibits squareroot cancellations within incomplete Kloosterman sums. It seems far beyond the current technique to prove Conjecture 8.1 for general $q$, even in the case $|I|\approx q^{0.99}.$ On the other hand, Conjecture 8.1 corresponds to the exponent pair hypothesis in the theory of arithmetic exponent pairs, for which we hope $(\kappa,\lambda)=(0,\frac{1}{2})$ is admissible for (4.2). This illustrates that all of our previous arguments should never beat (8.1) unconditionally even imposing factorization conditions on $q$.
However, as $\varpi\rightarrow 9/20-$ for instance, Theorem 1.5 shows that $C(\varpi)=5/(5-6\varpi)$ is admissible, as long as $q$ is squarefree and smooth. This is obviously sharper than the conjectural bound 8.1 which takes $C(\varpi)=6/(5-6\varpi).$ As one may see from previous sections, Theorem 1.5 is proved by using orthogonality of multiplicative characters, so that Lemma 4.6 on character sums has been applied to finish the proof. This phenomenon indicates that for $\varpi<1/2-$ it is clever to use multiplicative characters, and for $\varpi>1/2+$ additive ones might be more powerful instead. The exact transition range is very mysterious to the authors.